### v1

# Optimally generating $\mathfrak{su}(2^{N})$ using Pauli strings

Abstract
Any quantum computation consists of a sequence of unitary evolutions described by a finite set of Hamiltonians. When this set is taken to consist of only products of Pauli operators, we show that the minimal such set generating $\mathfrak{su}(2^{N})$ contains $12N+$ 1 elements. We provide a number of examples of such generating sets and furthermore provide an algorithm for producing a sequence of rotations corresponding to any given Pauli rotation, which is shown to have optimal complexity.
preprint: APS/123-QED
I Introduction
Performing a quantum computation requires the control of a quantum system. The study of the controllability of quantum systems has a substantial history with developments across both mathematics and physics [1, 2, 3, 4, 5, 6, 7, 8, 9]. The techniques typically applied in control theory are often drawn from algebra [10], with much focus on Lie algebra in the quantum context [1, 8].
In the context of quantum computing, control refers to the evolution of a closed quantum system from a known initial state to a state representing the logical output of the computation [11, 12]. As the evolution of a closed quantum system is described by an element of the (special) unitary group, the corresponding Lie algebra (and the subalgebras thereof), whose elements generate unitary evolution, have received much interest. Due to their relevance for system dynamics, these Lie algebras are often called dynamical Lie algebras (DLAs) [8, 13, 14]. With the current rise of quantum machine learning and parameterized quantum circuits, DLAs have received renewed attention with connections to phenomena such as over-parametrisation and barren plateaus being established [15, 16, 17].
DLAs, as with any Lie algebra, are vector spaces equipped with an associated operation called a Lie bracket. In the context of unitary evolution of a (finite) quantum system, the vector space is taken to be a vector space of matrices and the Lie bracket is taken to be the matrix commutator. In typical scenarios, the DLA of a quantum system is generated from a smaller set $\mathcal{A}$ of “elementary operations” using linear combinations and nested commutators. If this DLA describes all elements of the special unitary group, then $\mathcal{A}$ is said to be a universal generating set. In the context of quantum computing, $\mathcal{A}$ corresponds to the elementary gates, while the corresponding DLA describes the set of unitaries that can be implemented by composing these gates. Significant past work has aimed to classify and characterize such universal sets for a variety of quantum systems [12, 7, 18, 8, 4, 9].
Here, we continue in this vein with a specific focus on minimality. That is, we aim to understand how small a generating set $\mathcal{A}$ can be, while still maintaining universality. If no further restriction is made on $\mathcal{A}$, then the answer is already known: the minimal generating set for $\mathfrak{su}(2^{N})$ has only $2$ elements (see Theorem 6 in Ref. [3]). However, the elements of this set consist of large sums of Pauli operators and are not particularly practical from a quantum information point of view. For quantum computing purposes, a more natural generating set would be constructed from tensor products of Pauli operators (Pauli strings), as these arise as native gate sets of certain quantum architectures (see e.g., [19]). While the Pauli strings generated by physical interactions are typically restricted to single-body local fields or two-body Ising-like terms, quantum gates corresponding to non-local Pauli strings can also emerge naturally in certain schemes of measurement-based quantum computation [20, 21, 22, 23, 24, 25] as well as architectures based on quantum cellular automata [26, 27]. In such cases, determining the DLA and understanding how to efficiently generate a given element of the DLA are non-trivial problems that must be solved in order to make full use of the available quantum operations.
In this paper, we demonstrate that, for generating sets consisting solely of Pauli strings, the minimal number of operators required to generate $\mathfrak{su}(2^{N})$ is $12N+$ 1. The proof is constructive and is separated into two components. In 1, we give a construction for producing sets that generate $\mathfrak{su}(2^{N})$ from sets that are defined for smaller numbers of qubits. As a corollary of this result and from an explicit example of a Pauli generating set of $\mathfrak{su}(4)$, we are able to conclude that generating sets for $\mathfrak{su}(2^{N})$ of size $12N+$ 1 exist. In 2, we complete the proof that $12N+$ 1 is indeed the minimal size of any such generating set by showing any set of size $2N$ cannot generate the entire Lie algebra.
Alongside the proof of minimality, we provide an algorithm (Algorithm 1) that takes a generating set of Pauli strings and a target Pauli string as input, and outputs the sequence of generators that producing the target Pauli string via nested commutation. As described in further detail in a later section, this sequence can be related to the sequence of control unitaries that can produce the rotation of the target Pauli string through any angle. The algorithm is demonstrably optimal as it produces sequences of length $\mathcal{O}(N)$ and may be of interest for use in circuit compilation for architectures hosting native gate sets of Pauli strings.
The remainder of this manuscript is structured as follows. In Section II we provide a brief discourse regarding certain relevant aspects of quantum control theory and establish the notation and definitions pertinent to our results. In Section III, we highlight certain aspects unique to considering generating sets of products of Paulis, before presenting our results and a number of examples. Section IV concludes and the appendices contain the proofs of the results presented in the main text.
II Background
Let $\mathcal{H}$ denote a finite-dimensional Hilbert space and let |$\ket{\psi}\in\mathcal{H}$ be a state. The evolution of |$\ket{\psi}$ is governed by the Schrödinger equation
$i\hbar\frac{d}{dt}\ket{\psi}=H\ket{\psi}$ (1)
where $H$ is a Hermitian operator. Now suppose we have a specific set of such Hermitian operators $\{H_{i}\}$ under consideration. A natural question to ask is then: for two states |$\ket{\psi_{0}},\ket{\psi_{1}}$ is there some element of $\{H_{i}\}$ and some time $t_{1}$ such that |$\ket{\psi_{0}}$ evolves to |$\ket{\psi_{1}}$ at time $t_{1}$ via (1)?
In the context of quantum control theory, the set of operators $\{H_{i}\}$ represent the different ways in which the evolution of the quantum system can be controlled. If different control Hamiltonians can be applied at different times, one can ask a more refined version of the above question: is there a sequence of elements of $\{H_{i}\}$, denoted $H_{l_{1}},...,H_{l_{m}}$, and a sequence of times $t_{l_{1}},...,t_{l_{m}}$ such that evolving |$\ket{\psi_{0}}$ first by $H_{l_{1}}$ for time $t_{l_{1}}$ and each subsequent $H_{l_{j}}$ for time $t_{l_{j}}$ ultimately produces |$\ket{\psi_{1}}$? Taking each $H_{i}$ to be time-independent, this question is the same as asking whether
|$\ket{\psi_{1}}=e^{-iH_{l_{m}}t_{l_{j}}}\dots e^{-iH_{l_{1}}t_{l_{1}}}\ket{\psi_{0}}$ (2)
holds for some $H_{l_{1}},...,H_{l_{m}}\in\{H_{i}\}$ and $t_{l_{1}},...,t_{l_{m}}\in\mathbb{R}$ (note we are taking $\hbar=1$).
Each factor in the product on the right-hand side of (2) is a unitary operator. The correspondence between Hermitian operators and unitary ones is one example of the correspondence between Lie groups and their associated Lie algebras. For the purposes of this work, a (matrix) Lie group is a closed subgroup $G$ of the group of $k\times k$ invertible complex matrices, denoted $GL(k)$, for some finite $k$. For example, the set of $2^{N}\times 2^{N}$ unitary matrices with unit determinant, the so-called special unitary matrices denoted $SU(2^{N})$, is a Lie group and the most relevant one for the remainder of this work. A Lie algebra $\mathfrak{g}$ is a real subspace of the space of all $k\times k$ complex matrices $M_{k}(\mathbb{C})$ for some $k$, along with a bilinear map $[\cdot,\cdot]:\mathfrak{g}\times\mathfrak{g}\rightarrow\mathfrak{g}$ called the Lie bracket. The Lie bracket is required to satisfy the following properties for all $A,B,C\in\mathfrak{g}$:
  1. 1.
    $[A,B]=-[B,A]$;
  2. 2.
    $[A,[B,C]]+[B,[C,A]]+[C,[A,B]]=0$.
As we only consider matrix Lie groups, we take the Lie bracket to be based on the matrix commutator, i.e., for all $A,B\in M_{k}(\mathbb{C})$:
$[A,B]:=AB-BA.$ (3)
Furthermore, it will be convenient below to consider the family of maps, called adjoint endomorphisms, defined for each $A\in\mathfrak{g}$ via
$\begin{gathered}ad_{A}:\mathfrak{g}\rightarrow\mathfrak{g}\\ B\mapsto[A,B].\end{gathered}$ (4)
As each Lie group is associated to a Lie algebra, it is possible to work almost solely with the latter and it is often much more convenient to do so. Furthermore, this aligns more closely with the statement of the questions raised earlier, i.e., about sequences of operators $iH_{l}t_{l}$ rather than about the associated unitaries $e^{-iH_{l}t_{l}}$. In what follows, we consider the Lie algebra associated to $SU(2^{N})$, denoted by $\mathfrak{su}(2^{N})$ and defined to be the vector subspace of $M_{2^{N}}(\mathbb{C})$ given by the set
$\{A\in M_{2^{N}}(\mathbb{C})|\operatorname{Tr}[A]=0,A^{\dagger}=-A\}$ (5)
equipped with the matrix commutator as the Lie bracket. The reader is directed to e.g., Ref. [28] for a more thorough treatment of Lie groups and Lie algebras.
Consider again (2) and the associated question; can the Lie algebra formalism help us to answer it? The affirmative answer follows from the judicious use of the Baker-Campbell-Hausdorff (BCH) formula: for any Lie algebra $\mathfrak{g}$ and $A,B\in\mathfrak{g}$,
$e^{A}e^{B}=e^{A+B+\sum_{s=1}^{\infty}\frac{1}{s!}ad_{A}^{(s)}(B)}$ (6)
where $ad_{A}^{(s)}(B)$ denotes the map $ad_{A}$ composed with itself $s$ times and applied to $B$. Taking $A=-iH_{l_{2}}t_{l_{2}}$ and $B=-iH_{l_{1}}t_{l_{1}}$ we see that we can rewrite the right-most terms of the product in (2) as a real linear combination of $-iH_{l_{1}}$, $-iH_{l_{2}}$, their commutator, the commutator of $-iH_{l_{2}}$ with their commutator, and so on. By repeated use of the BCH formula, the entire product of (2) can be reduced to a single term whose exponent is a linear combination of the $H_{l_{j}}$ and nested commutators thereof.
We have thus demonstrated that a question of sequences of evolutions involving the $iH_{l_{j}}$ separately can be reduced to a single evolution involving combinations of the $iH_{l_{j}}$ and their nested commutators. If our aim is to determine whether there exists a sequence such that (2) holds for any pair of $N$-qubit states |$\ket{\psi_{0}},\ket{\psi_{1}}$ then it suffices to show that the $iH_{l_{j}}$ and their nested commutators generate $\mathfrak{su}(2^{N})$. Taking $\mathcal{A}:=\{iH_{i}\}$ to be a set of generators, the dynamical Lie algebra generated by $\mathcal{A}$ is
$\langle\mathcal{A}\rangle_{\mathrm{Lie}}:=\mathrm{span}_{\mathbb{R}}\left\{\mathcal{A}\bigcup_{r=1}^{\infty}\mathcal{A}_{ad^{(r)}}\right\}$ (7)
where
$\mathcal{A}_{ad^{(r)}}:=\{ad_{A_{i_{1}}}\cdots ad_{A_{i_{r}}}(A)|(A_{i_{1}},...,A_{i_{r}},A)\in\mathcal{A}^{r+1}\}.$ (8)
The question of controllability between any two states for a given set of generators $\mathcal{A}$ is thus a question of whether
$\langle\mathcal{A}\rangle_{\mathrm{Lie}}=\mathfrak{su}(2^{N}).$ (9)
Like any vector space, $\mathfrak{su}(2^{N})$ has a basis. For example, for $N=1$, we have that $\mathfrak{su}(2)=\textrm{span}_{\mathbb{R}}\{iX,iZ,iY\}$ where $X,Y,Z$ are the Hermitian $2\times 2$ Pauli matrices. Extending to any $N$, let us consider the set of all Pauli strings on $N$ qubits, $\mathcal{P}_{N}:=\{I,X,Y,Z\}^{\otimes N}$, such that for every element $P\in\mathcal{P}_{N}$,
$P=P_{1}\otimes P_{2}\otimes\cdots\otimes P_{N}$
where $P_{i}\in\{I,X,Y,Z\}$ acts on qubit $i=1,...,N$ and $I$ is the identity matrix. $\mathcal{P}_{N}$ is a basis (over $\mathbb{R}$) of the vector space of $2^{N}\times 2^{N}$ Hermitian matrices. Defining $\mathcal{P}_{N}^{*}:=\mathcal{P}_{N}-\{I^{\otimes N}\}$, we see that $i\mathcal{P}_{N}^{*}$ forms a basis of the $2^{N}\times 2^{N}$ traceless, skew-Hermitian matrices and hence $\mathfrak{su}(2^{N})=\textrm{span}_{\mathbb{R}}(i\mathcal{P}_{N}^{*})$. Defining
$\langle\mathcal{A}\rangle_{[\cdot,\cdot]}:=\mathcal{A}\bigcup_{r=1}^{\infty}\mathcal{A}_{ad^{(r)}}$ (10)
it follows that a sufficient condition for (9) to hold for a given set $\mathcal{A}$, is that for each $A\in i\mathcal{P}_{N}^{*}$ there is some non-zero $\alpha\in\mathbb{R}$ such that $\alpha A\in\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$. In such a case, we say that $\mathcal{A}$ is adjoint universal for $i\mathcal{P}_{N}^{*}$.
III Results
In this paper, we consider the question: what size is the smallest set $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$ such that $\langle\mathcal{A}\rangle_{\mathrm{Lie}}=\mathfrak{su}(2^{N})$? As noted in the introduction, if we don’t make the restriction that $\mathcal{A}$ be a subset of $i\mathcal{P}_{N}^{*}$ but allow $\mathcal{A}$ to be any set of traceless, skew-Hermitian matrices, then the answer to the analogous question is known to be $2$ [3].
The restriction to considering subsets of $i\mathcal{P}_{N}^{*}$ comes with a number of convenient features worth highlighting. First, for any elements $P,Q$ of $\mathcal{P}_{N}$, the matrix commutator of $P$ and $Q$ takes a particularly simple form:
$[P,Q]=\begin{cases}0&\>\>\mathrm{if}\>\>P,Q\>\>\mathrm{commute,}\\ 2PQ&\>\>\mathrm{if}\>\>P,Q\>\>\mathrm{anticommute.}\end{cases}$ (11)
In fact, for all $P,Q\in\mathcal{P}_{N}$ such that $[P,Q]\neq 0$, we have that $PQ=\pm 2iR$ for some $R\in\mathcal{P}_{N}$. Accordingly, the Lie bracket of $\mathfrak{su}(2^{N})$ takes elements of $i\mathcal{P}_{N}^{*}$ to elements of $i\mathcal{P}_{N}^{*}$ up to some real scalar. Pursuant to the discussion at the end of the previous section, when working with generating sets $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$, $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}=i\mathcal{P}_{N}^{*}$ is in fact also a necessary condition for $\langle\mathcal{A}\rangle_{\textrm{Lie}}=\mathfrak{su}(2^{N})$. This allows us to work at the level of $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ rather than the full DLA $\langle\mathcal{A}\rangle_{\textrm{Lie}}$.
A second feature of working with $i\mathcal{P}_{N}^{*}$ relates to the BCH formula (recall Equation 6). Using the above fact that strings of Pauli’s either commute or anticommute, it is possible to derive the following. For $A,B\in i\mathcal{P}_{N}^{*}$ such that $[A,B]\neq 0$, we have
$e^{\frac{\pi}{4}A}e^{tB}e^{-\frac{\pi}{4}A}=e^{\frac{t}{2}[A,B]}.$ (12)
The details of this calculation are given in Appendix A. By iteratively applying the above expression, we see that the evolution described by any element of $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ for a given $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$ can be realised via a sequence of evolutions of the same fixed duration (except one). Below, we consider such sets $\mathcal{A}$ for which the elements of $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ are proportional to $i\mathcal{P}_{N}^{*}$ and so in such a case the above expression produces a method for performing any Pauli rotation in depth linear in $N$ as we will see in Section III.2.
Finally, as $i\mathcal{P}_{N}^{*}$ is intimately related to the Pauli group on $N$ qubits, there is machinery developed for the latter that can be leveraged to make statements about the former. Specifically, there is a canonical mapping between the $N$-qubit Pauli group and the additive group $\mathbb{F}_{2}^{2N}$ ($\mathbb{F}_{2}$ denotes the additive group on two elements), that we make use of to prove certain results below (namely 2). For simplicity, we delay explication of this mapping until Appendix B, however it is commonly used in stabiliser quantum mechanics for example to efficiently simulate Clifford quantum circuits classically [29] and to demonstrate that graph states are a canonical form for stabiliser states [30].
In the remainder of this section, we demonstrate (cf. 1) that $\mathfrak{su}(2^{N})$ can be generated by some $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$ such that $|\mathcal{A}|=2N+1$. We establish this fact by first demonstrating that it is possible to construct a set generating $\mathfrak{su}(2^{N})$ from a set of generators of $\mathfrak{su}(2^{k})$ for some $2\leq k<N$ in conjunction with a set of $2(N-k)$ elements satisfying certain properties (cf. 1). The claim of $12N+$ 1 then follows from noting that $\mathfrak{su}(4)$ can be generated using $5$ elements from $i\mathcal{P}_{2}^{*}$ (cf 1).
To show that generating sets of size $12N+$ 1 are indeed optimal, we prove that any subset of $i\mathcal{P}_{N}^{*}$ consisting of $2N$ elements fails to generate $\mathfrak{su}(2^{N})$ (cf. 2). Since the maximal size of any subset $\mathcal{A}^{\prime}$ of $i\mathcal{P}_{N}^{*}$ containing independent elements is $2N$, it follows that any generating set $\mathcal{A}$ drawn from $i\mathcal{P}_{N}^{*}$ necessarily contains a dependent element.
Finally, we note that a key part of the proof of 1 consists in constructing sequences of elements of $\mathcal{A}$ that produce any given element of $i\mathcal{P}_{N}^{*}$ (up to some non-zero real scalar) using commutators. We provide an algorithm for producing such sequences (see Algorithm 1) which may be of interest in the context of circuit compilation for circuits using a gate set related to those in Theorem 1. We furthermore show that this algorithm returns sequences of size $\mathcal{O}(N)$ which is optimal (see Theorem 3).
III.1 Minimal generating sets of $\mathfrak{su}(2^{N})$
Before stating our first theorem, we require a notion similar to adjoint universality discussed earlier. We say that a set $\mathcal{B}\subset i\mathcal{P}_{N}^{*}$ is product universal if the set
×$\langle\mathcal{B}\rangle_{\times}:=\bigcup_{r=1}^{\infty}\{B_{i_{1}}\cdot B_{i_{2}}\cdot...\cdot B_{i_{r}}|(B_{i_{1}},...,B_{i_{r}})\in\mathcal{B}^{r}\}$ (13)
is such that for each $A\in i\mathcal{P}_{N}^{*}$ there is an $\alpha\in\{\pm 1,\pm i\}$ such that $\alpha A\in\langle\mathcal{B}\rangle_{\times}$ (note the difference in the allowed set of scalars compared to adjoint universality). The statement of the theorem is then:
Theorem 1.
Fix $k\geq 2$ and consider $\mathcal{A}\subset i\mathcal{P}_{k}^{*}$ and $\mathcal{B}\subset i\mathcal{P}^{*}_{N-k}$ such that $\langle\mathcal{A}\rangle_{\mathrm{Lie}}=\mathfrak{su}(2^{k})$ and $\mathcal{B}$ is product universal for $i\mathcal{P}^{*}_{N-k}$. Let $\mathbb{B}$ be a set of pairs $\{(U_{B},B)\}_{B\in\mathcal{B}}$ where each $U_{B}\in\mathcal{P}_{k}^{*}$. Defining $\mathcal{A}^{\prime}=\{A\otimes I^{\otimes N-k}|A\in\mathcal{A}\}$ and =$\mathcal{B}^{\prime}=\{U_{B}\otimes B|(U_{B},B)\in\mathbb{B}\}$, we have that
$\langle\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}\rangle_{\mathrm{Lie}}=\mathfrak{su}(2^{N}).$
The proof of this theorem is given in Appendix C. In simple terms, 1 shows that any set of operators that generates $\mathfrak{su}(2^{k})$ for $k\geq 2$ can be combined with a set of operators that generate a basis of $\mathfrak{su}(2^{N-k})$ under multiplication to form a generating set of $\mathfrak{su}(2^{N})$ by extending the latter non-trivially on the other $k$ qubits.
To arrive at our claim that there exist generating subsets of $i\mathcal{P}_{N}^{*}$ consisting of $12N+$ 1 elements, we also require the following result:
Proposition 1.
For $\mathcal{A}=\{iX_{1},iZ_{1},iX_{2},iZ_{2},iZ_{1}\otimes Z_{2}\}$, we have that
$\langle\mathcal{A}\rangle_{\textrm{Lie}}=\mathfrak{su}(4).$ (14)
The proof of the proposition is a direct calculation that $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}=i\mathcal{P}_{2}^{*}$, which is provided in Appendix D. Noting that there exist product universal sets for $i\mathcal{P}_{m}^{*}$ consisting of $2m$ elements, we thus have the following:
Corollary 1.
For $\mathcal{A}$ as in 1, and $\mathcal{A}^{\prime}$ and $\mathcal{B}^{\prime}$ as in 1 with $|\mathcal{B}^{\prime}|=(2N-2)$, $|\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}|=2N+1$.
Having shown that it is possible to generate $\mathfrak{su}(2^{N})$ from a set $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$ such that $|\mathcal{A}|=2N+1$, to show that this is optimal requires a proof that no strictly smaller set suffices. In fact, we need only check that no generating set $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$ of $2N$ elements is able to produce the entire Lie algebra, for the following reasons. For any $\mathcal{A},\mathcal{A}^{\prime}\subset i\mathcal{P}_{N}^{*}$ such that $\mathcal{A}\subseteq\mathcal{A}^{\prime}$, we have that
$\langle\mathcal{A}\rangle_{[\cdot,\cdot]}\subseteq\langle\mathcal{A}^{\prime}\rangle_{[\cdot,\cdot]}$ (15)
and so a proof that any set of size $2N$ is insufficient immediately implies that any set contained within it is insufficient also. The following theorem shows that it is indeed impossible to generate $\mathfrak{su}(2^{N})$ using $2N$ elements of $i\mathcal{P}_{N}^{*}$:
Theorem 2.
Let $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$ consist of $2N$ elements. Then,
$\langle\mathcal{A}\rangle_{\mathrm{Lie}}\neq\mathfrak{su}(2^{N}).$
The proof is given in Appendix I and consists in showing that there exists an element $A$ of $i\mathcal{P}_{N}^{*}$ such that $\alpha A\notin\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ for any non-zero $\alpha\in\mathbb{R}$. Note that we can assume that $\mathcal{A}$ contains only independent elements since if this were not the case then there exists an element of $i\mathcal{P}_{N}^{*}$ independent of all elements of $\mathcal{A}$ which is thus necessarily not an element of $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$. With this assumption, we consider two cases: first, where $\mathcal{A}$ consists only of elements that all mutually anticommute, and second, where $\mathcal{A}$ contains at least two distinct elements that commute. For the first case, it is readily shown that for three distinct elements $A_{1},A_{2},A_{3}\in\mathcal{A}$, no element proportional to $A_{1}A_{2}A_{3}$ is in $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ (cf. 1 for a formal proof). For the second case, we show that for any $A_{1},A_{2}\in\mathcal{A}$ that commute, no element proportional to $A_{1}A_{2}$ is in $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$, by showing that no sequence of elements of $\mathcal{A}$ of even length (cf. 3) nor odd length (cf. 2) produces the necessary element by nested commutators. The proof of this theorem makes use of the connection between the Pauli group on $N$ qubits and $\mathbb{F}_{2}^{2N}$, including a symplectic structure on the latter, which is explained in detail in Appendix B.
To summarise the results so far, we have seen that to generate $\mathfrak{su}(2^{N})$ using elements of $i\mathcal{P}_{N}^{*}$, one needs at least $12N+$ 1 elements including an element not independent of the others. Two particularly interesting examples of such minimal generating sets are the following.
Example 1: Consider the following sets for 1:
$\mathcal{A}$ =$=\{iX_{1},iZ_{1},iX_{2},iZ_{2},iZ_{1}\otimes Z_{2}\}$
$\mathcal{B}^{\prime}_{1}$ =$=\{iX_{2}\otimes\prod_{2<j<i}Y_{j}\otimes Z_{i},iZ_{2}\otimes\prod_{2<j<i}Y_{j}\otimes X_{i}\}_{i=3,...,N}.$ (16)
These sets do obviously fulfill the conditions of Theorem 1 (given Proposition 1) and therefore give an adjoint universal set of size $12N+$ 1. This set may not be particularly interesting in itself, but becomes more relevant in light of the next example.
Example 2: Consider the following sets for 1:
$\mathcal{A}$ =$=\{iX_{1},iZ_{1},iX_{2},iZ_{2},iZ_{1}\otimes Z_{2}\}$
$\mathcal{B}^{\prime}_{2}$ =$=\{iX_{i}\otimes Z_{i+1},iZ_{i}\otimes X_{i+1}\}_{i=2,...,N-1}.$ (17)
These sets are particularly interesting as they contain only nearest neighbor interactions and have $12N+$ 1 elements. However, they do not fulfill the conditions of Theorem 1 because not all elements of $\mathcal{B}^{\prime}_{2}$ have support on qubits 1 and 2. Fortunately, we can show $\mathcal{B}^{\prime}_{1}\subset\langle\mathcal{B}_{2}^{\prime}\rangle_{[\cdot,\cdot]}$ (up to some factors) as follows:
$iX_{2}\otimes\prod_{2<j<i}Y_{j}\otimes Z_{i}$ $\propto ad_{X_{2}Z_{3}}\cdots ad_{X_{i-2}Z_{i-1}}(X_{i-1}Z_{i})$
$iZ_{2}\otimes\prod_{2<j<i}Y_{j}\otimes X_{i}$ $\propto ad_{Z_{2}X_{3}}\cdots ad_{Z_{i-2}X_{i-1}}(Z_{i-1}X_{i})$
Since $\mathcal{A}^{\prime}\cup\mathcal{B}_{1}^{\prime}$ is adjoint universal, so is $\mathcal{A}^{\prime}\cup\mathcal{B}_{2}^{\prime}$. From this we see that Theorem 1 does not only give us a way to construct optimal generating set, but also allows us to prove universality for others.
In Algorithm 1, we introduce the PauliCompiler which is an optimal compiler for minimal adjoint universal sets of operators such as Eq. 16. But it presupposes sets as in Theorem 1. That is, if we intend to use it to compile the set $\mathcal{A}^{\prime}\cup\mathcal{B}_{2}^{\prime}$ from Eq. 17, we first need to map each element as shown above. Since this preprocessing requires $\mathcal{O}(N)$ in the average case and PauliCompiler nests the input generators, the algorithm will produce sequences of length $\mathcal{O}(N^{2})$.
III.2 Optimal compiler for 1
Algorithm 1 PauliCompiler$(V\otimes W)$
Target: $V\otimes W\in i\mathcal{P}^{*}_{N}$.
            $\mathcal{A},\mathcal{A}^{\prime},\mathcal{B},\mathcal{B}^{\prime}$ as in Theorem 1.
if $W=I$ then
      Choose $A_{1},...,A_{s}\in\mathcal{A}$ s.t. $V\propto ad_{A_{1}}\cdots ad_{A_{s-1}}(A_{s})$ $\wedge\>s=\mathcal{O}(1)$
return $(A_{1}\otimes I,...,A_{s}\otimes I)$
else if $V=I$ then
     Choose $W_{1},W_{2}\in i\mathcal{P}_{N-k}$ s.t. $W\propto[W_{1},W_{2}]$
      $\mathcal{G}^{\prime}=(G_{1}^{\prime},...,G_{\mathcal{G}^{\prime}}^{\prime})\leftarrow\textsc{SubsystemCompiler}(W_{1})$ s.t. $V_{1}^{\prime}\otimes W_{1}\propto ad_{G^{\prime}_{|\mathcal{G}^{\prime}|}}\cdots ad_{G_{2}^{\prime}}(G_{1}^{\prime})$
      $\mathcal{G}^{\prime\prime}=(G_{1}^{\prime\prime},...,G_{\mathcal{G}^{\prime\prime}}^{\prime\prime})\leftarrow\textsc{SubsystemCompiler}(W_{2})$ s.t. $V_{2}^{\prime}\otimes W_{2}\propto ad_{G^{\prime\prime}_{|\mathcal{G}^{\prime\prime}|}}\cdots ad_{G_{2}^{\prime\prime}}(G_{1}^{\prime\prime})$
      Choose $A_{1},...,A_{s}\in\mathcal{A}$ s.t. $V_{1}^{\prime}\propto ad_{A_{1}}\cdots ad_{A_{s}}(V_{2}^{\prime})$ $\wedge\>s=\mathcal{O}(1)$
     $\mathcal{G}=(G_{1},...,G_{|\mathcal{G}|})\equiv\mathcal{G}^{\prime}\oplus\mathcal{G}^{\prime\prime}\oplus(A_{1}\otimes I,...,A_{s}\otimes I)$
     Choose $\sigma\in S_{|\mathcal{G}|}$ s.t. $ad_{G_{\sigma(1)}}\cdots ad_{G_{\sigma(|\mathcal{G}|-1)}}(G_{\sigma(|\mathcal{G}|)})\neq 0$
return $(G_{\sigma(1)},...,G_{\sigma(|\mathcal{G}|)})$
else
      $\mathcal{G}^{\prime}=(G_{1}^{\prime},...,G_{\mathcal{G}^{\prime}}^{\prime})\leftarrow\textsc{SubsystemCompiler}(W)$ s.t. $V^{\prime}\otimes W\propto ad_{G^{\prime}_{|\mathcal{G}^{\prime}|}}\cdots ad_{G_{2}^{\prime}}(G_{1}^{\prime})$
      Choose $A_{1},...,A_{s}\in\mathcal{A}$ s.t. $V\propto ad_{A_{1}}\cdots ad_{A_{s}}(V^{\prime})$ $\wedge\>s=\mathcal{O}(1)$
return $({A_{1}\otimes I,...,A_{s}\otimes I,G^{\prime}_{|\mathcal{G}^{\prime}|},...,G^{\prime}_{1}})$
end if
Algorithm 2 SubsystemCompiler(W)
Target: $W\in i\mathcal{P}^{*}_{N-k}$.
            $\mathcal{A},\mathcal{A}^{\prime},\mathcal{B},\mathcal{B}^{\prime}$ as in Theorem 1.
$\textrm{Choose }U_{1}\otimes B_{1},...,U_{r}\otimes B_{r}\in\mathcal{B}^{\prime}\textrm{ s.t. }\prod_{i=1,...,r}B_{i}\propto W$ $\wedge\>r=\mathcal{O}(N)$
$i\leftarrow r-1$
$\mathcal{G}=(U_{r}\otimes B_{r},\>)$
$\mathcal{H}=\emptyset$
while $i\geq 1$ do
     if $\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A\propto I$ then
[…]
         $\mathcal{H}\leftarrow(A_{1},A_{2})$
         $\mathcal{G}\textrm{.append}(A_{1}\otimes I,A_{2}\otimes I)$
     else if $\left[U_{i}\otimes B_{i},\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I\right]=0$ then
[…]
         $\mathcal{H}\leftarrow A^{\prime}$
         $\mathcal{G}.\textrm{append}(A^{\prime}\otimes I)$
     else
         $\mathcal{G}.\textrm{append}(U_{i}\otimes B_{i})$
         $i=i-1$
     end if
end while
return $\mathcal{G}=(G_{1},...,G_{|\mathcal{G}|})\textrm{ s.t. }\exists\alpha\in\mathbb{R}^{*}\wedge V^{\prime}\in\mathcal{P}_{k}^{*}\textrm{ with }$ $V^{\prime}\otimes W=\alpha\>ad_{G_{|\mathcal{G}|}}\cdots ad_{G_{2}}(G_{1})$
Let us turn our attention to the question of how to optimally compile any Pauli operator $P\in\langle\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}\rangle_{[\cdot,\cdot]}$. That is, we want to find a sequence of nested commutators of generating Pauli strings that equals a desired target Pauli string. A simple brute-force method to do this involves exploring all possible sequences of commutators in a tree-like manner until all Pauli strings are found, which will give sequences of optimal length for each string. However, since there is no way to target a specific Pauli string, this process will take exponentially long in the number of qubits in the worst case. In contrast, our algorithm takes polynomial time in the number of qubits for any desired Pauli string.
We show the following Theorem:
Theorem 3.
Consider a set of operators $\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ as in Theorem 1 such that $|\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}|=2N+1$. Let $V\otimes W\in i\mathcal{P}_{N}$. Then, PauliCompiler($V\otimes W$) of Algorithm 1 returns a sequence of operators $G_{1},...,G_{L}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ such that $V\otimes W=\alpha\>ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})$ for some $\alpha\in\mathbb{R}^{*}$ and $L=\mathcal{O}(N)$.
The full proof of this Theorem is given in Appendix J in combination with Appendices F and G where we prove that both Algorithms 1 and 2 terminate and are correct.
The central component of PauliCompiler in Algorithm 1 is the SubsystemCompiler of Algorithm 2 which uses a specific choice of “helper” Paulis in $\mathcal{A}$ such that any product $W\propto\prod_{i}B_{i}$ for $B_{i}\in\mathcal{B}$ can be generated under Lie brackets as $V^{\prime}\otimes W$ for some $V^{\prime}\in\mathcal{P}^{*}_{k}$ (see Appendix G). However, this does not give us control over $V^{\prime}$. Therefore, we state the following proposition:
Proposition 2.
Let $V,V^{\prime}\in i\mathcal{P}^{*}_{k}$ for any $k\geq 1$ and $\mathcal{A}$ as in Theorem 1. Then, there exist $A_{1},...,A_{r}\in\mathcal{A}$ and $\alpha\in\mathbb{R}^{*}$ such that $V=\alpha\>ad_{A_{1}}\cdots ad_{A_{r}}(V^{\prime})$.
This proposition is proven in Appendix H. Using this proposition, we can then always compile an operator $V\otimes W$ for $V\in\mathcal{P}^{*}_{k}$ and $W\in\mathcal{P}^{*}_{N-1}$ (which is the ELSE statement in Algorithm 1). The other IF statements in Algorithm 1 cover the cases $W=I$ and $V=I$. The former is trivial due to the adjoint universality of $\mathcal{A}$. The latter is more complicated as we need to create two operators $V\otimes W_{1},V\otimes W_{2}\in i\mathcal{P}_{N}$ such that $[V\otimes W_{1},V\otimes W_{2}]\propto I\otimes W$. Fortunately, we can create those operators because of the SubsystemCompiler and Proposition 2. However, their commutator is not a nested commutator but a commutator of commutators. Fortunately, we can show in Appendix H, that we can always find a re-ordering of operators such that a commutator of commutators can be written as a nested commutator (up to a sign).
The proof that the returned sequence of PauliCompiler has length $L=\mathcal{O}(N)$, is given in Appendix J and follows directly from the construction of the algorithms. Note that $L=\mathcal{O}(N)$ is optimal for a set $\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ of size $12N+$ 1. To see this, consider a product universal set $\mathcal{B}\subset\mathcal{P}_{N-k}$ as in Theorem 1 that is minimal, i.e., $|\mathcal{B}|=2(N-k)$. Let $P=\prod_{B\in\mathcal{B}}B$. Then, there exist no other product $P\propto\prod_{i=1,...,L}B_{i}$ with $L<2(N-k)$. At the same time, every other $P^{\prime}\in\mathcal{P}_{N-k}$ can be written as $P^{\prime}\propto\prod_{i=1,...,L}B_{i}$ for some $L<2(N-k)$. This is a direct consequence of the formalism we discuss in Appendix B. That means that $\mathcal{O}(N)$ is optimal for product universality. Since $k$ is a constant and the commutator is strictly weaker than the product (see Eq. 11)), $L=\mathcal{O}(N)$, as it can be achieved by PauliCompiler, must also be optimal.
Let us conclude by showing how the results presented in this paper allow us to optimally implement any Pauli rotation given access to Hamiltonians of the form $\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ as in Theorem 1 with $|\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}|=2N+1$. Consider a sequence $G_{1},...,G_{L}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ returned by PauliCompiler. Given Eq. 12, we implement an arbitrary rotation around the product of this sequence as follows:
$\exp{\left(\frac{t}{2^{L-1}}ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})\right)}$
$=e^{\frac{\pi}{4}G_{1}}\cdots e^{\frac{\pi}{4}G_{L-1}}e^{tG_{L}}e^{-\frac{\pi}{4}G_{L-1}}\cdots e^{-\frac{\pi}{4}G_{1}}$ (18)
Since $L=\mathcal{O}(N)$, so is the length of this sequence, which is optimal for a minimal generating set as discussed above.
IV Discussion
As quantum computation and quantum control are so intimately connected, any pursuit of a fully-fledged quantum computer benefits from both experimental and theoretical progress made in the latter field. In this work, we have contributed to this progress by deepening the understanding of which sets of control Hamiltonians are required for universal quantum control.
Specifically, we demonstrated that when the set of control Hamiltonians is taken to consist entirely of products of Pauli operators on $N$ qubits, a minimum of at least $12N+$ 1 different Hamiltonians are required. We provide a simple prescription on how to construct such minimal sets, which does not only provide interesting control Hamiltonians but can also be used as a theoretical tool to prove universality of other control Hamiltonians. Additionally, we give an algorithm that, when given such as set of control Hamiltonians $\mathcal{A}$ and a specific Pauli string as input, outputs a sequence of elements of $\mathcal{A}$ describing an implementation of any rotation of the input Pauli string. This algorithm has polynomial runtime in the number of qubits and is optimal in the sense that it produces sequences of minimal length. Our algorithm is tailored to generating sets of a specific form, namely the form described in Theorem 1, but we believe its principles can be applied more generally.
There are a number of reasons why understanding quantum control involving products of Paulis is of relevance. First, two-qubit Pauli rotations are native in certain quantum computing architectures such as in trapped ions [19]. Consequently, the results presented here are directly applicable to questions of control of such systems. Second, in the context of quantum combinatorial optimisation, typically the clauses of the constraint function under consideration are mapped to strings of Pauli-$Z$ operators (see e.g., [31, 32]). It is conceivable that a better understanding of the universality of circuits comprised of product of Pauli rotations can lead to improvements in the design of quantum optimisation algorithms. Finally, for quantum computational schemes that are not based on universal few-body gates (the “standard gate set”) but instead involve gates generated by long Pauli strings [21, 27, 25], our results provide crucial insights for assessing universality and solving the compilation problem.
During the preparation of this manuscript, we became aware of a related work, Ref. [33], which contains some similar results to our own. However, their results are reached via a different methods to that which we employ here. It would be an interesting avenue for future work to determine whether the different approaches could be combined to make further progress in this area.
Acknowledgements.
The authors thank Vedran Dunjko, Paul Barth and Arunava Majumder for stimulating discussions. This project was funded in whole or in part by the Austrian Science Fund (FWF) [DK-ALM W1259125912591259-N27272727, SFB BeyondC F7102, WIT9503323]. For open access purposes, the authors have applied a CC BY public copyright license to any author-accepted manuscript version arising from this submission. This work was also co-funded by the European Union (ERC, QuantAI, Project No. $101055129$). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them. DTS is supported by the Simons Collaboration on Ultra-Quantum Matter, which is a grant from the Simons Foundation (651440).
Appendix A BCH Formula for Paulis
Here, we spell out the details regarding the reduction of the BCH formula for elements of $i\mathcal{P}_{N}^{*}$ expressed in Equation 12 in the main text.
Let $A,B\in i\mathcal{P}_{N}^{*}$ be such that $[A,B]\neq 0$, i.e. that $[A,B]=2AB$ or equivalently $AB=-BA$. It follows that
$[A,[A,B]]=2[A,AB]=2(A^{2}B-ABA)=4A^{2}B=-4B$ (19)
where the last equality uses the fact that elements of $i\mathcal{P}_{N}^{2}$ square to minus the identity. Using the BCH formula, we have
$e^{\frac{\pi}{4}A}e^{tB}e^{-\frac{\pi}{4}A}=e^{tB+\sum_{k=1}^{\infty}\frac{1}{k!}ad_{\frac{\pi}{4}A}^{(k)}(tB)}$ (20)
where $ad_{\frac{\pi}{4}A}^{(k)}(tB)$ denotes the composition of $ad_{\frac{\pi}{4}A}$ $k$-times and applied to $tB$, as in the main text. Using the above commutation relations, we see that
$tB+\sum_{k=1}^{\infty}\frac{1}{k!}ad_{\frac{\pi}{4}A}^{(k)}(tB)$ $=t\left\{B+\sum_{k=1}^{\infty}\frac{1}{k!}(\frac{\pi}{4})^{k}ad_{A}^{(k)}(B)\right\}$ (21)
$=t\left\{B\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}}{(2k)!}(\frac{\pi}{4})^{2k}2^{2k}\right)+[A,B]\left(\sum_{k=0}^{\infty}\frac{(-1)^{k}}{(2k+1)!}(\frac{\pi}{4})^{2k}2^{2k}\right)\right\}$ (22)
$=t\left(\cos(\frac{\pi}{2})B+\frac{1}{2}\sin(\frac{\pi}{2})[A,B]\right)$ (23)
$=\frac{t}{2}[A,B].$ (24)
Appendix B The Pauli Group and $\mathbb{F}_{2}^{2N}$
In this appendix, we explain the connection between the Pauli group on $N$ qubits and the symplectic space $\mathbb{F}_{2}^{2N}$ ($\mathbb{F}_{2}$ denotes the finite field containing two elements) alluded to in the main text and used in a number of proofs below.
Let us denote by $\mathcal{G}_{\mathcal{P}_{N}}$ the Pauli group on $N$ qubits defined via
$\mathcal{G}_{\mathcal{P}_{N}}:=\{e^{i\theta}P|\theta\in\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\},P\in\mathcal{P}_{N}\}$ (25)
with group multiplication given by the standard matrix multiplication. As in the main text, $\mathcal{P}_{N}$ denotes the set of Pauli strings. Since the single qubit Pauli operators satisfy $Y=iXZ$, it is possible to write any element $e^{i\theta}P\in\mathcal{G}_{\mathcal{P}_{N}}$ as $e^{i\theta^{\prime}}P_{X}P_{Z}$ where $P_{X}\in\mathcal{P}_{N}$ is the Pauli string consisting of $X$ terms at every tensor factor where $P$ contained either an $X$ or a $Y$ and identities elsewhere, $P_{Z}\in\mathcal{P}_{N}$ is similarly defined with $Z$ terms at every factor where $P$ contained either a $Z$ or $Y$ and identities elsewhere, and $\theta^{\prime}=\theta+y\frac{\pi}{2}$ where $y$ is the number of $Y$ terms contained in $P$. We say that elements of $\mathcal{G}_{\mathcal{P}_{N}}$ written in this way are in standard form.
Let $e^{i\theta_{P}}P_{X}P_{Z}$ and $e^{i\theta_{Q}}Q_{X}Q_{Z}$ be elements of $\mathcal{G}_{\mathcal{P}_{N}}$ in standard form. We see that the standard from of their product can be written is
$\left(e^{i\theta_{P}}P_{X}P_{Z}\right)\left(e^{i\theta_{Q}}Q_{X}Q_{Z}\right)=e^{i(\theta_{P}+\theta_{Q}+\lambda_{PQ}\pi)}\left(P_{X}Q_{X}\right)\left(P_{Z}Q_{Z}\right)$ (26)
where $\lambda_{PQ}\in\{0,1\}$ is the parity of the number of tensor factors where $P_{Z}$ and $Q_{Z}$ both contain a non-identity term. Accordingly, the term $\lambda_{PQ}$ contains all the information about whether $e^{i\theta_{P}}P_{X}P_{Z}$ and $e^{i\theta_{Q}}Q_{X}Q_{Z}$ commute or not: $\lambda_{PQ}=0$ if they commute and $\lambda_{PQ}=1$ otherwise.
Clearly, $\mathcal{P}_{N}$ is a subset of $\mathcal{G}_{\mathcal{P}_{N}}$. When $\mathcal{S}\subset\mathcal{P}_{N}$ is such that (i) $\mathcal{S}$ is a subgroup of $\mathcal{G}_{\mathcal{P}_{N}}$, (ii) $-I\notin\mathcal{S}$ and (iii) all elements of $\mathcal{S}$ commute, then $\mathcal{S}$ is called a stabiliser subgroup [34] and is used throughout many areas of quantum computing and information. Many results arising in e.g., quantum error correction make use of the number of independent generator of the group $\mathcal{S}$, that is, the smallest number of elements of $\mathcal{S}$ that generate $\mathcal{S}$ and are not products of each other. A convenient way to check independence is via the map $f:\mathcal{S}\rightarrow\mathbb{F}_{2}^{2N}$ defined via
$e^{i\theta}P_{X}P_{Z}\mapsto\bm{v}_{P}$ (27)
where $e^{i\theta}P_{X}P_{Z}$ is the standard form of an element of $\mathcal{S}$ and $\bm{v}_{P}$ is the vector containing a $1$ at entry $j$ if the $j$th tensor factor of $P_{X}$ contains an $X$, a $1$ at entry $N+j$ if the $j$th tensor factor of $P_{Z}$ contains a $Z$, and zeros otherwise. It follows that the above map is a group isomorphism where $\mathbb{F}_{2}^{2N}$ is taken as an additive group, meaning that the product of elements in $\mathcal{S}$ corresponds to the sum of elements in $\mathbb{F}_{2}^{2N}$. In fact, it is convenient (and harmless) to take $\mathbb{F}_{2}^{2N}$ to be a vector space over $\mathbb{F}_{2}$. By doing so, the independence of elements of $\mathcal{S}$ can be rephrased as linear independence of the corresponding elements in $\mathbb{F}_{2}^{2N}$.
Let us now consider the map $f$ extended to all of $\mathcal{G}_{\mathcal{P}_{N}}$. This map remains a a group homomorphism however it is now many-to-one since it sends every element of $\mathcal{G}_{\mathcal{P}_{N}}$ with standard forms that differ only in the angle $\theta$ to the same element of $\mathbb{F}_{2}^{2N}$. Nevertheless, when restricted to $\mathcal{P}_{N}\subset\mathcal{G}_{\mathcal{P}_{N}}$ or $i\mathcal{P}_{N}^{*}$, this map remains useful for our purposes as many of the results presented in the main text make statements regarding equivalence up to proportionality. We make this intuition more precise below.
Before doing so, let us add one last piece of structure to $\mathbb{F}_{2}^{2N}$. Define the matrix $\Lambda\in M_{2N}(\mathbb{F}_{2})$ to be
$\Lambda=\begin{bmatrix}\bm{0}&I_{N}\\ I_{N}&\bm{0}\end{bmatrix}$ (28)
where $\bm{0}\in M_{N}(\mathbb{F}_{2})$ is the matrix of all zeros and $I_{N}\in M_{N}(\mathbb{F}_{2})$ is the identity matrix. Abusing notation slightly, we see that the symplectic bilinear form $\Lambda:\mathbb{F}_{2}^{2N}\times\mathbb{F}_{2}^{2N}\rightarrow\mathbb{F}_{2}$ defined via $(\bm{v},\bm{w})\mapsto\bm{v}^{\top}\Lambda\bm{w}$ turns the vector space $\mathbb{F}_{2}^{2N}$ into a symplectic vector space. This symplectic form captures the commutativity of elements of $\mathcal{G}_{\mathcal{P}_{N}}$ exactly: $P,Q\in\mathcal{G}_{\mathcal{P}_{N}}$ commute if and only if $\bm{v}:=f(P)$ and $\bm{w}:=f(Q)$ are such that $\bm{v}^{\top}\Lambda\bm{w}=0$.
The following lemma establishes the justification for being able to work with $\mathbb{F}_{2}^{2N}$ for proving certain statements made in the main text. For any subset $\mathcal{C}\subset\mathcal{G}_{\mathcal{P}_{N}}$, let $f(\mathcal{C})\subset\mathbb{F}_{2}^{2N}$ denote the image of $\mathcal{C}$. Note that, this in particular means that $f(i\mathcal{P}_{N}^{*})=\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$ where $\bm{0}\in\mathbb{F}_{2}^{2N}$ is the vector containing only 00 entries. For $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$, let us define
$\langle\mathcal{A}\rangle_{\|\cdot\|_{2}}:=\left\{\frac{A}{\|A\|_{2}}|A\in\langle\mathcal{A}\rangle_{[\cdot,\cdot]}\right\}$ (29)
where $\|\cdot\|_{2}$ denotes the spectral norm. Note that $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}^{\|\cdot\|_{2}}\subset i\mathcal{P}_{N}^{*}$. We have:
Lemma 1.
Let $\mathcal{A},\mathcal{B}\subset i\mathcal{P}_{N}^{*}$. Then
  1. (i)
    $\mathcal{A}$ is adjoint universal for $i\mathcal{P}_{N}^{*}$ if and only if $f(\langle\mathcal{A}\rangle_{\|\cdot\|_{2}})$ contains $\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$.
  2. (ii)
    $\mathcal{B}$ is product universal for $i\mathcal{P}_{N}^{*}$ if and only if $f(\mathcal{B})$ contains a basis of $\mathbb{F}_{2}^{2N}$.
Proof.
The proof essentially follows from the observation that there is a one-to-one correspondence between $i\mathcal{P}_{N}^{*}$ and $f(i\mathcal{P}_{N}^{*})=\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$, and hence also between any subset $\mathcal{C}\subset i\mathcal{P}_{N}^{*}$ and $f(\mathcal{C})$.
Case (i): Suppose that $\mathcal{A}$ is adjoint universal for $i\mathcal{P}_{N}^{*}$. This means that for each $A\in i\mathcal{P}_{N}^{*}$ there is some non-zero $\alpha\in\mathbb{R}$ such that $\alpha A\in\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$. In particular, this means that $\frac{\alpha}{|\alpha|}A\in\langle\mathcal{A}\rangle_{\|\cdot\|_{2}}$ and hence that $f(A)\in f(\langle\mathcal{A}\rangle_{\|\cdot\|_{2}})$. It follows that $f(i\mathcal{P}_{N}^{*})\subseteq f(\langle\mathcal{A}\rangle_{\|\cdot\|_{2}})$.
Suppose now that $f(\langle\mathcal{A}\rangle_{\|\cdot\|_{2}})$ contains $\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$. In particular, this means that $f(\langle\mathcal{A}\rangle_{\|\cdot\|_{2}})=\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$. Since each $\bm{v}\in\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$ corresponds to some $A\in i\mathcal{P}_{N}^{*}$, it follows that there is some element $A^{\prime}\in\langle\mathcal{A}\rangle_{\|\cdot\|_{2}}$ such that $A=e^{i\theta}A^{\prime}$ for some $\theta\in\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\}$. However, since both $A$ and $A^{\prime}$ are in $i\mathcal{P}_{N}^{*}$, it must be the case that $A=\pm A^{\prime}$. Since $A^{\prime}\in\langle\mathcal{A}\rangle_{\|\cdot\|_{2}}$ means that there is some $A^{\prime\prime}\in\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ such that $A^{\prime\prime}/\|A^{\prime\prime}\|_{2}=A^{\prime}$, it follows that $\alpha A\in\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ for $\alpha:=\|A^{\prime\prime}\|_{2}$.
Case (ii): Suppose that $\mathcal{B}$ is product universal for $i\mathcal{P}_{N}^{*}$. Recall that this means that for any $A\in i\mathcal{P}_{N}^{*}$, there exists a sequence $B_{i_{1}},B_{i_{2}},...,B_{i_{r}}\in\mathcal{B}$ such that $A=e^{i\theta}B_{i_{1}}...B_{i_{r}}$ for some $\theta\in\{0,\frac{\pi}{2},\pi,\frac{3\pi}{2}\}$. In particular, this means that $f(A)=f(B_{i_{1}}...B_{i_{r}})=f(B_{i_{1}})+...+f(B_{i_{r}})$. Noting that $f(A)\neq\bm{0}$ since $A\in i\mathcal{P}_{N}^{*}$ and hence also that $f(B_{i_{1}})+...+f(B_{i_{r}})\neq\bm{0}$. The latter implies that some subset of $\{f(B_{i_{1}}),...,f(B_{i_{r}})\}$ is linear independent. The union of all such subsets for each $A\in i\mathcal{P}_{N}^{*}$ necessarily contains a basis of $\mathbb{F}_{2}^{2N}$ since $f(i\mathcal{P}_{N}^{*})=\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$.
Suppose that $f(\mathcal{B})$ contains a basis of $\mathbb{F}_{2}^{2N}$. Let this basis be denoted by $\{\bm{v}_{1},...,\bm{v}_{2N}\}$. Since there is a one-to-one correspondence between $\mathcal{B}$ and $f(\mathcal{B})$, we can associate each $\bm{v}_{j}$ with some $B_{j}\in\mathcal{B}$. Since any element $\bm{w}\in\mathbb{F}_{2}^{2N}\setminus\{\bm{0}\}$ an be written as $\bm{w}=\bm{v}_{i_{1}}+...+\bm{v}_{i_{r}}=f(B_{i_{1}})+...+f(B_{i_{r}})=f(B_{i_{1}}...B_{i_{r}})$ and since $\bm{w}=f(A)$ for some $A\in i\mathcal{P}_{N}^{*}$, it follows that $A=e^{i\theta}B_{i_{1}}...B_{i_{r}}$ for some $\theta\in\{0,\frac{\pi}{2},\pi,\frac{3\pi}{4}\}$. ∎
Having demonstrated that it suffices to work entirely with $\mathbb{F}_{2}^{2N}$ for the purposes of this paper, let us introduce the following notation for later use. In direct parallel to Equation 4, let us define the map $ad_{\bm{v}}:\mathbb{F}_{2}^{2N}\rightarrow\mathbb{F}_{2}^{2N}$ via $ad_{\bm{v}}(\bm{w})=\left(\bm{v}^{\top}\Lambda\bm{w}\right)(\bm{v}+\bm{w})$. We note that, for all $A,B\in i\mathcal{P}_{N}^{*}$ we have that
$ad_{A}(B)=\begin{cases}\pm 2AB,&\text{ if }A,B\text{ anticommute},\\ 0&\text{ otherwise}\end{cases}$ (30)
and
$ad_{f(A)}(f(B))=\begin{cases}f(A)+f(B),&\text{ if }A,B\text{ anticommute },\\ 0&\text{ otherwise}.\end{cases}$ (31)
That is, $ad_{A}(B)\neq 0$ if and only if $ad_{f(A)}(f(B))\neq 0$, and moreover if $ad_{A}(B)\neq 0$, then
$f\left(\frac{ad_{A}(B)}{\|ad_{A}(B)\|_{2}}\right)=ad_{f(A)}(f(B)).$ (32)
Since, for any $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$,
$\textrm{span}_{\mathbb{R}}\langle\mathcal{A}\rangle_{[\cdot,\cdot]}=\textrm{span}_{\mathbb{R}}\langle\mathcal{A}\rangle_{\|\cdot\|_{2}},$ (33)
we will typically make no distinction between results adjoint universality proven in the context of $i\mathcal{P}_{N}^{*}$ and those proven in $\mathbb{F}_{2}^{2N}$, and drop reference to $\|\cdot\|_{2}$ from future discourse.
Appendix C Proof of Theorem 1
For convenience, we restate the Theorem 1 here:
Theorem 1.
Fix $k\geq 2$ and consider $\mathcal{A}\subset i\mathcal{P}_{k}^{*}$ and $\mathcal{B}\subset i\mathcal{P}^{*}_{N-k}$ such that $\langle\mathcal{A}\rangle_{\mathrm{Lie}}=\mathfrak{su}(2^{k})$ and $\mathcal{B}$ is product universal for $i\mathcal{P}^{*}_{N-k}$. Let $\mathbb{B}$ be a set of pairs $\{(U_{B},B)\}_{B\in\mathcal{B}}$ where each $U_{B}\in\mathcal{P}_{k}^{*}$. Defining $\mathcal{A}^{\prime}=\{A\otimes I^{\otimes N-k}|A\in\mathcal{A}\}$ and =$\mathcal{B}^{\prime}=\{U_{B}\otimes B|(U_{B},B)\in\mathbb{B}\}$, we have that
$\langle\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}\rangle_{\mathrm{Lie}}=\mathfrak{su}(2^{N}).$
Proof.
To prove Theorem 1, we prove that Algorithm 1 finishes and is correct. This is proven within the following Lemma:
Lemma 1.
For any input $V\otimes W\in i\mathcal{P}^{*}_{N}$ and sets $\mathcal{A}^{\prime},\mathcal{B}^{\prime}$ as in Theorem 1, Algorithm 1 finishes and returns a finite, ordered set of operators $(G_{1},...,G_{L})$ with $G_{i}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ for all $i=1,...,L$ such that $\exists\alpha\in\mathbb{R}^{*}$ with $ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})=\alpha\>V\otimes W$.
Proof.
The proof can be found in Appendix F. ∎
For a high-level discussion of Algorithm 1, we refer to Appendix E. Since the algorithm finishes for any input $V\otimes W\in i\mathcal{P}^{*}_{N}$, this means that $\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ is adjoint universal and therefore, $\textrm{span}_{\mathbb{R}}\langle\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}\rangle_{[\cdot,\cdot]}=\mathfrak{su}(2^{N})$. ∎
Appendix D Proof of 1
Proof of 1.
The proof proceeds by direct calculation. Note that we have
$i\mathcal{P}_{2}^{*}=\{iA\otimes B|A,B\in\{I,X,Y,Z\}\}\setminus\{iI\}.$ (34)
Since $[iX,iZ]=-2iY$, $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ clearly contains all elements of $i\mathcal{P}_{2}^{*}$ with a single non-identity term. By noting that
$[iX_{1},iZ_{1}\otimes Z_{2}]=-2iY_{1}\otimes Z_{2},$ (35)
$[iZ_{1},[iX_{1},iZ_{1}\otimes Z_{2}]=4iX_{1}\otimes Z_{2},$ (36)
and similarly for taking commutators with $X_{2}$ and $Z_{2}$, we see that $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$ contains all Pauli strings with two non-identity terms and hence $\langle\mathcal{A}\rangle_{[\cdot,\cdot]}=i\mathcal{P}_{2}^{*}$, proving the claim. ∎
Appendix E Details of Algorithm 1
For convenience, let us analyze Algorithm 1 in a bit more detail. The algorithm has been reprinted in Fig. 1 with line numbering. It relies on sets of operators $\mathcal{A}^{\prime},\mathcal{B}^{\prime}\subset i\mathcal{P}_{N}$ constructed from $\mathcal{A}\in i\mathcal{P}^{*}_{k}$ and $\mathcal{B}\in i\mathcal{P}_{N-k}$ as defined in Theorem 1. Their particular structure is crucial for the workings of the algorithm. In particular, while $\mathcal{B}$ is only product universal, each element of $\mathcal{B}^{\prime}$ must also act on the first $k$-qubit subspace. Only because of that can we make use of the adjoint universality of $\mathcal{A}$ to promote $\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ to adjoint universality on the full space using SubsystemCompiler of Algorithm 2 which is reprinted in Fig. 2.
Given the input $V\otimes W\in i\mathcal{P}^{*}_{N}$, PauliCompiler distinguishes three cases:
Case 1: $W=I$ (lines 2-4). Here, we make use of the fact that $\mathcal{A}$ is adjoint universal to produce any operator $V\otimes I$ for $V\in i\mathcal{P}_{k}$.
Case 2: $W\neq I$ and $V\neq I$ (lines 13-16). In line 14, we rely on SubsystemCompiler$(W)$ to compile a sequence $V^{\prime}\otimes W$ for some $V^{\prime}\in\mathcal{P}^{*}_{k}$. Given Propostion 2, we can then map this to an arbitrary $V$ (up to a factor).
Case 3: $V=I$ (lines 5-12). This is the most complicated case. Instead of generating $I\otimes W$ directly, we compile two $V_{1}^{\prime}\otimes W_{1}$ and $V_{1}^{\prime}\otimes W_{2}$ (using the same procedure as in case 2) such that $[V_{1}^{\prime}\otimes W_{1},V_{1}^{\prime}\otimes W_{2}]\propto I\otimes W$. However, it is not obvious that for two operators $B=ad_{B_{1}}\cdots ad_{B_{r-1}}(B_{r})$ and $C=ad_{C_{1}}\cdots ad_{C_{l-1}}(C_{l})$, we can write $[B,C]=ad_{D_{1}}\cdots ad_{D_{m-1}}(D_{m})$. As we prove in Lemma 2, we can reorder the elements (as in line 11) such that the commutator of commutators is a nested commutator.
To implement the algorithm, we write a table for $\mathcal{A}$ that lists for each $V\in i\mathcal{P}_{k}$ a sequence $A_{1},...,A_{s}$ as in line 3. Then, one writes another table that lists for each pair $V,V^{\prime}\in\mathcal{P}_{k}$ a sequence $A_{1},...,A_{s}$ as in lines 9 and 15. To generate the nested commutator from the commutators of commutators in line 11 (and implicitly in lines 9 and 15 due to Proposition 2), one iteratively applies Lemma 2. SubsystemCompiler can be implemented directly as described in the pseudocode.
Appendix F Proof of Lemma 1
Algorithm 3 PauliCompiler$(V\otimes W)$
1:Target: $V\otimes W\in i\mathcal{P}^{*}_{N}$.
2:            $\mathcal{A},\mathcal{A}^{\prime},\mathcal{B},\mathcal{B}^{\prime}$ as in Theorem 1.
3:if $W=I$ then
4:     Choose $A_{1},...,A_{s}\in\mathcal{A}$ s.t. $V\propto ad_{A_{1}}\cdots ad_{A_{s-1}}(A_{s})\wedge s=\mathcal{O}(1)$$\triangleright$ Use adjoint universality of $\mathcal{A}$.
5:return $(A_{1}\otimes I,...,A_{s}\otimes I)$
6:else if $V=I$ then
7:     Choose $W_{1},W_{2}\in i\mathcal{P}_{N-k}$ s.t. $W\propto[W_{1},W_{2}]$
8:      $\mathcal{G}^{\prime}=(G_{1}^{\prime},...,G_{\mathcal{G}^{\prime}}^{\prime})\leftarrow\textsc{SubsystemCompiler}(W_{1})$ s.t. $V_{1}^{\prime}\otimes W_{1}\propto ad_{G^{\prime}_{|\mathcal{G}^{\prime}|}}\cdots ad_{G_{2}^{\prime}}(G_{1}^{\prime})$triangleright Construct $V_{1}^{\prime}\otimes W_{1}$ for a given $W_{1}$ and some $V_{1}^{\prime}\in\mathcal{P}^{*}_{k}$
9:      $\mathcal{G}^{\prime\prime}=(G_{1}^{\prime\prime},...,G_{\mathcal{G}^{\prime\prime}}^{\prime\prime})\leftarrow\textsc{SubsystemCompiler}(W_{2})$ s.t. $V_{2}^{\prime}\otimes W_{2}\propto ad_{G^{\prime\prime}_{|\mathcal{G}^{\prime\prime}|}}\cdots ad_{G_{2}^{\prime\prime}}(G_{1}^{\prime\prime})$triangleright Construct $V_{2}^{\prime}\otimes W_{2}$ for a given $W_{2}$ and some $V_{2}^{\prime}\in\mathcal{P}^{*}_{k}$
10:     Choose $A_{1},...,A_{s}\in\mathcal{A}$ s.t. $V_{1}^{\prime}\propto ad_{A_{1}}\cdots ad_{A_{s}}(V_{2}^{\prime})\wedge s=\mathcal{O}(1)$triangleright Use Proposition 2.
11:     $\mathcal{G}=(G_{1},...,G_{|\mathcal{G}|})\equiv\mathcal{G}^{\prime}\oplus\mathcal{G}^{\prime\prime}\oplus(A_{1}\otimes I,...,A_{s}\otimes I)$$\triangleright$ Construct $V_{1}^{\prime}\otimes W_{1},V_{1}^{\prime}\otimes W_{2}$ s.t. $[V_{1}^{\prime}\otimes W_{1},V_{1}^{\prime}\otimes W_{2}]\propto I\otimes W$.
12:     Choose $\sigma\in S_{|\mathcal{G}|}$ s.t. $ad_{G_{\sigma(1)}}\cdots ad_{G_{\sigma(|\mathcal{G}|-1)}}(G_{\sigma(|\mathcal{G}|)})\neq 0$$\triangleright$ Use Lemma 2
13:return $(G_{\sigma(1)},...,G_{\sigma(|\mathcal{G}|)})$
14:else
15:      $\mathcal{G}^{\prime}=(G_{1}^{\prime},...,G_{\mathcal{G}^{\prime}}^{\prime})\leftarrow\textsc{SubsystemCompiler}(W)$ s.t. $V^{\prime}\otimes W\propto ad_{G^{\prime}_{|\mathcal{G}^{\prime}|}}\cdots ad_{G_{2}^{\prime}}(G_{1}^{\prime})$triangleright Construct $V^{\prime}\otimes W$ for a given $W$ and some $V\in\mathcal{P}^{*}_{k}$
16:     Choose $A_{1},...,A_{s}\in\mathcal{A}$ s.t. $V\propto ad_{A_{1}}\cdots ad_{A_{s}}(V^{\prime})\wedge s=\mathcal{O}(1)$$\triangleright$ Use Proposition 2.
17:return $({A_{1}\otimes I,...,A_{s}\otimes I,G^{\prime}_{|\mathcal{G}^{\prime}|},...,G^{\prime}_{1}})$
18:end if
Figure 1: Algorithm 1 from the main text with additional line numbering and comments. Here, we write $\propto$ to indicate proportionality but nonzero and $A\oplus B$ is the concatenation of two sequences $A,B$.
For a high-level discussion of Algorithm 1, we refer to Appendix E. For convenience, we restate Lemma 1 here:
Lemma 1.
For any input $V\otimes W\in i\mathcal{P}^{*}_{N}$ and sets $\mathcal{A}^{\prime},\mathcal{B}^{\prime}$ as in Theorem 1, Algorithm 1 finishes and returns a finite, ordered set of operators $(G_{1},...,G_{L})$ with $G_{i}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ for all $i=1,...,L$ such that $\exists\alpha\in\mathbb{R}^{*}$ with $ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})=\alpha\>V\otimes W$.
Proof.
The algorithm is reprinted in Fig. 1 with additional line numbering which will be referred to in this proof. As this algorithm relies on the SubsystemCompiler of Algorithm 2, we refer to Lemma 1 in Sec. G to prove that SubsystemCompiler is correct and finishes. Here, we rely on it to return a sequence of operators whose nested commutator is proportional to $V^{\prime}\otimes W$ given $W\in i\mathcal{P}^{*}_{N-k}$ for some $V\in i\mathcal{P}^{*}_{k}$.
Statement 1: The algorithm is correct. We show that the algorithm is correct for any input through the four following statements.
Statement 1.1: All choices in the algorithm are valid. There are five choices in the algorithms, in lines 3, 6, 9, 11 and 15.
Line 3: Since by construction $\mathcal{A}\subset i\mathcal{P}_{k}$ is adjoint universal, for each $V\in i\mathcal{P}_{k}$ there exist a finite sequence $A_{1},...,A_{s}\in\mathcal{A}$ such that $V\propto ad_{A_{1}}\cdots ad_{A_{s-1}}(A_{s})$ is nonzero. $s=\mathcal{O}(f(k))$ for some function $f$ of $k$. Since $k$ is constant, so is $s$.
Line 6. For any Pauli operator $W\in\mathcal{P}_{N}$ there exist two other Pauli operators $W_{1},W_{2}\in\mathcal{P}_{N}$ that anticommute and $W_{1}W_{2}\propto W$. This is trivially true for $N=1$ and easily extends to all other $N$ by induction.
Line 9 and 15. Proven by Propostion 2 in Appendix H. From the proof of Proposition 2, we can immediately infer that the chosen sequences may be finite and have length $s=\mathcal{O}(f(k))$ for some function $f$ of $k$. Since $k$ is constant, so is $s$.
Line 11. By construction, we have
$I\otimes W$ $\propto[V_{1}^{\prime}\otimes W_{1},V_{1}^{\prime}\otimes W_{2}]$
$\propto[V_{1}^{\prime}\otimes W_{1},ad_{A_{1}}\cdots ad_{A_{s}}(V_{2}^{\prime})\otimes W_{2}]$
$\propto[ad_{G^{\prime}_{|\mathcal{G}^{\prime}|}}\cdots ad_{G_{2}^{\prime}}(G_{1}^{\prime}),ad_{A_{1}\otimes I}\cdots ad_{A_{s}\otimes I}ad_{G^{\prime\prime}_{|\mathcal{G}^{\prime\prime}|}}\cdots ad_{G_{2}^{\prime\prime}}(G_{1}^{\prime\prime})]$ (37)
$=ad_{ad_{G^{\prime}_{|\mathcal{G}^{\prime}|}}\cdots ad_{G_{2}^{\prime}}(G_{1}^{\prime})}\left(ad_{A_{1}\otimes I}\cdots ad_{A_{s}\otimes I}ad_{G^{\prime\prime}_{|\mathcal{G}^{\prime\prime}|}}\cdots ad_{G_{2}^{\prime\prime}}(G_{1}^{\prime\prime})\right)\neq 0$
This is a commutator of commutators instead of the nested commutator. As we prove in Lemma 2 in Appendix H, any nonzero commutator of commutators can be written as a nonzero nested commutator with reordered elements.
Statement 1.2: The returned sequence $G_{1},...,G_{L}$ is such that $G_{1},...,G_{L}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$. This is true by construction of the algorithm.
Statement 1.3: The returned sequence $G_{1},...,G_{L}$ is such that $ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})\neq 0$. This is true by construction of the algorithm.
Statement 1.4: The returned sequence $G_{1},...,G_{L}$ is such that $ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})\propto V\otimes W$. This is true by construction of the algorithm.
Statement 2: The algorithm finishes. Lemma 1 shows that SubsystemCompiler finishes and returns a finite sequence. Since all choices are valid and return finite sequences, the algorithm must finish. ∎
Appendix G Proof of Algorithm 2
Algorithm 4 SubsystemCompiler($W$)
1:Target: $W\in i\mathcal{P}^{*}_{N-k}$. $\mathcal{A},\mathcal{A}^{\prime},\mathcal{B},\mathcal{B}^{\prime}$ as in Theorem 1.
2:$\textrm{Choose }U_{1}\otimes B_{1},...,U_{r}\otimes B_{r}\in\mathcal{B}^{\prime}\textrm{ s.t. }\prod_{i=1,...,r}B_{i}\propto W\wedge r=\mathcal{O}(N)$$\triangleright$ Use product universality of $\mathcal{B}$.
3:$i\leftarrow r-1$
[…]
$\mathcal{H}=\emptyset$
6:while $i\geq 1$ do
7:     if $\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A\propto I$ then $\triangleright$ Use “helper” Paulis to ensure that no identities appear in the product.
[…]
9:         $\mathcal{H}\leftarrow(A_{1},A_{2})$
10:         $\mathcal{G}\textrm{.append}(A_{1}\otimes I,A_{2}\otimes I)$triangleright Add helper Paulis to sequence.
11:     else if $\left[U_{i}\otimes B_{i},\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I\right]=0$ then $\triangleright$ Use helper Pauli to make commuting elements anticommute.
[…]
13:         $\mathcal{H}\leftarrow A^{\prime}$
14:         $\mathcal{G}.\textrm{append}(A^{\prime}\otimes I)$triangleright Add helper Pauli to sequence.
15:     else
16:         $\mathcal{G}.\textrm{append}(U_{i}\otimes B_{i})$triangleright Add next element of the product to the sequence.
17:         $i=i-1$
18:     end if
19:end while
20:return $\mathcal{G}=(G_{1},...,G_{|\mathcal{G}|})$triangleright $[G_{|\mathcal{G}|},[\cdots,[G_{2},G_{1}]\cdots]]=\alpha V^{\prime}\otimes W$ for $V^{\prime}\neq I$ and $\alpha\in\mathbb{R}^{*}$
Figure 2: Algorithm 2 from the main text with additional line numbering and comments.
As part of the proof of Lemma 1 in Appendix F, let us prove that Algorithm 2 is correct and finishes through the following Lemma:
Lemma 1.
For any input $W\in i\mathcal{P}^{*}_{N-k}$, Algorithm 2 finishes and returns a finite, ordered set of operators $\mathcal{G}=(G_{1},...,G_{|\mathcal{G}|})$ with $G_{i}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ for all $i=1,...,|\mathcal{G}|$ such that $\exists\alpha\in\mathbb{R}^{*}$ with $ad_{G_{|\mathcal{G}|}}\cdots ad_{G_{2}}(G_{1})=\alpha V^{\prime}\otimes W$ for some $V^{\prime}\in\mathcal{P}^{*}_{k}$.
Proof.
The algorithm is reprinted in Fig. 2 with additional line numbering which will be referred to in this proof.
Statement 1: The algorithm is correct. Let us start by showing that the algorithm is correct, i.e., it returns a sequence of operators $G_{1},...,G_{|\mathcal{G}|}$ such that $ad_{G_{|\mathcal{G}|}}\cdots ad_{G_{2}}(G_{1})\propto V^{\prime}\otimes W$ for some $V^{\prime}\in\mathcal{P}^{*}_{k}$ is nonzero. Given a target Pauli $W\in i\mathcal{P}_{N-k}$, in line 1, we first find a sequence of operators $U_{1}\otimes B_{1},...,U_{r}\otimes B_{r}\in\mathcal{B}^{\prime}$ such that $\prod_{i=1,...,r}B_{i}\propto W$. This is always guaranteed to exist since $\mathcal{B}$ is product universal. As we discuss in Appendix J, the optimal choice is $r=\mathcal{O}(N)$. However, it might be that $ad_{U_{1}\otimes B_{1}}\cdots ad_{U_{r-1}\otimes B_{r-1}}(U_{r}\otimes B_{r})=0$. Therefore, the algorithm goes iteratively through this sequence, starting with $r$ and adds operators $A\otimes I\in\mathcal{A}^{\prime}$ between the product to ensure that the nested commutator never gives zero. Let us prove the correctness through four further statements:
Statement 1.1: All choices in the algorithm are valid. There are three choices being made in the algorithm, in line 1, 7 and 11. We already proved the existence of the choice in line 1 with $r=\mathcal{O}(N)$.
Line 7: For simplicity, we identify $A\equiv A_{1}$, $B\equiv A_{2}$ and $C\equiv U_{i}$. Let $C\in\mathcal{P}^{*}_{k}$ be arbitrary. ($U_{i}\neq I$ by construction of $\mathcal{B}^{\prime}$.) Then, we show that there exist $A,B\in\mathcal{P}^{*}_{k}$ such that $A\neq B$ and $[A,C]\neq 0$ and $[B,C]\neq 0$ and $[A,B]=0$ iff $k>1$. This is trivially not true for $k=1$ and trivially true for $k=2$. The rest follows easily by induction. To see this, consider $C_{1}\otimes C_{2}\in\mathcal{P}^{*}_{k+1}$ for $C_{1}\in\mathcal{P}_{k}$ and $C_{2}\in\mathcal{P}_{1}$. If $C_{1}\neq I$ there exist such $A,B$ by assumption. If $C_{1}=I$, then choose $A=X\otimes X\otimes A_{2}$ and $B=Z\otimes Z\otimes A_{2}$ with $A_{2}\in\mathcal{P}_{1}$ such that $[A_{2},C]\neq 0$ (where we assumed an identity on all but the first two and last qubits).
Line 11: For simplicity, we identify $A\equiv A^{\prime}$, $B\equiv\prod_{j>i}U_{j}\prod_{A\in\mathcal{H}}A$ and $C\equiv U_{i}$. Let $B\neq C\in\mathcal{P}^{*}_{k}$ be arbitrary. (As we will see below, for any $i$, if we get to line 11, then $B\neq I$.) Then, we show that there exist $A\in\mathcal{P}^{*}_{k}$ such that $[A,B]\neq 0$ and $[A,C]\neq 0$ and $A\not\propto BC$ iff $k>1$. Note that for any two Pauli operators $B\neq C\in\mathcal{P}^{*}_{k}$, we can find another Pauli $A\in\mathcal{P}_{k}$ that anticommutes with both. However, for $k=1$ this Pauli must be the product of the other two. Consider $k>1$ for two commuting Pauli operators $B\neq C$. Since then, $[BC,B]=0$, any such anticommuting $A$ cannot be equal to $BC$. Now, consider two mutually anticommuting Pauli operators $B\neq C$. Choose $A^{\prime}=BC$ which anticommutes with both $B$ and $C$. Since $k>1$, there exist a Pauli operator $A^{\prime\prime}\in\mathcal{P}_{k}^{*}$ that commutes with both $B$ and $C$. Choose $A=A^{\prime}A^{\prime\prime}$.
Let us now see that the requirement $B\neq I$ is given, i.e., $\prod_{j>i}U_{j}\prod_{A\in\mathcal{H}}A\neq I$ at line 11 for any iteration $i$. Clearly, this is true for the first iteration $i=r-1$. Now note that at iteration $i$ of the algorithm, the next step can only be reached through line 16 where $i=i-1$. This line can only be accessed once lines 6 and 10 evaluate to false. This means that at line 16, $\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A\neq I$. Going to the next iteration $i=i-1$, at line 6 the previous inequality reads $\prod_{j>i}U_{j}\prod_{A\in\mathcal{H}}A\neq I$. If line 6 evaluates to false, this is also true at line 11. If line 6 evaluates to true, then the constraints on $A_{1},A_{2}$ are such that $A_{1}A_{2}\neq U_{i}$ and therefore $\prod_{j>i}U_{j}\prod_{A\in\mathcal{H}\cup\{A_{1},A_{2}\}}A\neq I$. This can be seen as follows: Assume $A_{1}A_{2}=U_{i}$. Then, $[A_{1},A_{2}]=0$ and therefore, $[A_{1},A_{1}A_{2}]=0$. However, $[A_{1},A_{1}A_{2}]=[A_{1},U_{1}]\neq 0$, which is a contradiction. That is, we can assume $B\neq I$ above.
Statement 1.2: $G_{1},...,G_{|\mathcal{G}|}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$. This is true by construction of the algorithm.
Statement 1.3: $ad_{G_{|\mathcal{G}|}}\cdots ad_{G_{2}}(G_{1})\neq 0$. We must show that for any $l=1,...,|\mathcal{G}|$, $[G_{l},\prod_{j<l}G_{j}]\neq 0$. There is two cases: $G_{l}\in\mathcal{B}^{\prime}$ and $G_{l}\in\mathcal{A}^{\prime}$.
Let $G_{l}\in\mathcal{B}^{\prime}$: This means that the algorithm has reached line 15 for some iteration. Therefore, we know that line 10 evaluated as false for some $G_{l}=U_{i}\otimes B_{i}$. At that iteration $i$ of the algorithm, $\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I=\alpha\prod_{j<l}G_{l}$ for some $\alpha\in\mathbb{R}^{*}$. Therefore, $[G_{l},\prod_{j<l}G_{l}]\neq 0$.
Let $G_{l}\in\mathcal{A}^{\prime}$. This operator has been added at some iteration $i$ in line 7 or 11 of the algorithm. Assume first that $G_{l}$ has been added at line 7 at iteration $i$. At that iteration, we have $\prod_{j<l}G_{j}=\alpha\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I$ for some $\alpha\in\mathbb{R}^{*}$. At line 7, we also have $U_{i}\propto\prod_{j>i}U_{j}\prod_{A\in\mathcal{H}}A$. $A_{1},A_{2}$ are both chosen such that they mututally commute, but both anticommute with $U_{i}$ and hence, in both cases, $[G_{l},\prod_{j<l}G_{j}]\neq 0$. Now assume $G_{l}$ has been added at line 11 at iteration $i$. At that iteration, we have $\prod_{j<l}G_{j}=\alpha\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I$ for some $\alpha\in\mathbb{R}^{*}$. $A^{\prime}$ is chosen such that it anticommutes with $\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I$ and therefore, $[G_{l},\prod_{j<l}G_{j}]\neq 0$.
Statement 1.4: $ad_{G_{|\mathcal{G}|}}\cdots ad_{G_{2}}(G_{1})\propto V^{\prime}\otimes W$. Since $ad_{U_{1}\otimes B_{1}}\cdots ad_{U_{r-1}\otimes B_{r-1}}(U_{r}\otimes B_{r})\propto V^{\prime\prime}\otimes W$ and the algorithm only adds operators $A\otimes I\in\mathcal{A}^{\prime}$, $ad_{G_{|\mathcal{G}|}}\cdots ad_{G_{2}}(G_{1})\propto V^{\prime}\otimes W$
Statement 2: The algorithm finishes. Let us conclude by showing that the algorithm finishes, i.e., it returns the output in finite steps. The algorithm would not finish if the while loop always evaluates as true, i.e. $i$ is never reduced to 0. $i$ is reduced only in line 16, so we have to make sure that for any $i$, both line 6 and 10 will eventually evaluate as false. Let us show this by proving the following statements:
Statement 2.1: If line 6 evaluates as true at some iteration, adding $A_{1},A_{2}$ (as chosen in line 7) to $\mathcal{H}$ will make it evaluate as false in the following loop. If at some iteration $i$, $\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A\propto I$, then $A_{1}A_{2}\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A\propto A_{1}A_{2}$. Since $A_{1}\neq A_{2}$ as by the choice in line 7, the statement is true.
Statement 2.2: If line 10 evaluates as true at some iteration, adding $A^{\prime}$ (as chosen in line 11) to $\mathcal{H}$ will make it evaluate as false in the following loop. This is due to the first constraint in line 11. If at some iteration $i$, $\left[U_{i}\otimes B_{i},\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I\right]=0$, then, $\left[U_{i}\otimes B_{i},A^{\prime}\otimes I\prod_{j>i}U_{j}\otimes B_{j}\prod_{A\in\mathcal{H}}A\otimes I\right]\neq 0$ because $[U_{i}\otimes B_{i},A_{i}\otimes I]\neq 0$. Hence, the statement is true.
Statement 2.3: If line 10 evaluates as true at some iteration, adding $A^{\prime}$ (as chosen in line 11) to $\mathcal{H}$ will make line 6 evaluate as false in the following loop. This is due to the third constraint in line 11. Since $A^{\prime}\not\propto\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A$ as by the choice in line 11, $A^{\prime}\prod_{j\geq i}U_{j}\prod_{A\in\mathcal{H}}A\not\propto I$ and therefore, line 6 evaluates as false.
Since all choices are provably valid and return finite sequences, line 16 will deterministically be reached and the algorithm finishes. This also implies that the returned sequence is finite. ∎
Appendix H Proof of Proposition 2
For convenience, we restate Propostion 2 here:
Proposition 1.
Let $V,V^{\prime}\in i\mathcal{P}^{*}_{k}$ for any $k\geq 1$ and $\mathcal{A}$ as in Theorem 1. Then, there exist $A_{1},...,A_{r}\in\mathcal{A}$ and $\alpha\in\mathbb{R}^{*}$ such that $V=\alpha\>ad_{A_{1}}\cdots ad_{A_{r}}(V^{\prime})$.
Proof.
The Proposition is proven through Lemmas 1 and 2 below. In the first Lemma, we prove that we can always find a sequence of Paulis that map one Pauli to another under commutation.
Lemma 1.
Let $V,V^{\prime}\in i\mathcal{P}_{k}^{*}$ for any $k\in\mathbb{N}^{*}$. Then there exist $P_{1},...,P_{r}\in i\mathcal{P}_{k}^{*}$ and $\alpha\in\mathbb{R}^{*}$ such that $V=\alpha\>ad_{P_{1}}\cdots ad_{P_{r}}(V^{\prime})$.
Proof.
We prove Lemma 1 by induction. $N=1$ is trivial. Now consider $V=iV_{1}\otimes V_{2}$ and $V^{\prime}=iV_{1}^{\prime}\otimes V_{2}^{\prime}$ for $V_{1},V_{1}^{\prime}\in\mathcal{P}_{k}$ and $V_{2},V_{2}^{\prime}\in\mathcal{P}_{1}$. We consider two cases.
Case 1: $V_{1},V_{2},V_{1}^{\prime},V_{2}^{\prime}\neq I$. By assumption, there exist $P_{1}\otimes I,...,P_{r}\otimes I$ such that $V_{1}\otimes V_{2}^{\prime}\propto ad_{P_{1}}\cdots ad_{P_{r}}(V_{1}^{\prime}\otimes V_{2}^{\prime})$. Analogous to the case $N=1$, $V^{\prime}_{2}$ can be mapped to an operator proportional to $V_{2}$.
Case 2: $V_{1}=I$ or $V_{2}=I$ or $V_{1}^{\prime}=I$ or $V_{2}^{\prime}=I$. If $V_{1}^{\prime}=I$, then $V_{2}^{\prime}\neq I$ and vice verse. If $V_{1}^{\prime}=I$, then choose $P_{r}=A\otimes B$ with $A\neq I$ and $B\neq V_{2}^{\prime}$ such that $[P_{r},I\otimes V_{2}^{\prime}]=\alpha\>A\otimes V_{2}^{\prime}B$ for $\alpha\in\mathbb{R}^{*}$. If $V_{2}^{\prime}=I$, then choose $P_{r}=A\otimes B$ with $A\neq V_{1}^{\prime}$ and $B\neq I$ such that $[P_{r},V_{1}^{\prime}\otimes I]=\alpha\>AV_{1}^{\prime}\otimes B$ for $\alpha\in\mathbb{R}^{*}$. If $V_{1}\neq I$ and $V_{2}\neq I$, then this reduces to case 1. So from hereon, w.l.o.g. assume that $V^{\prime}_{1},V_{2}^{\prime}\neq I$.
If $V_{1}=I$, then $V_{2}\neq I$ and vice verse. If $V_{1}=I$, then consider case 1 for a target $V=A\otimes V_{2}B$ such that $A\neq I$ and $[V_{2},B]\neq 0$. Then, choose $P_{0}=A\otimes B$. If $V_{2}=I$, then consider case 1 for a target $V=V_{1}A\otimes B$ such that $[V_{1},A]\neq 0$ and $B\neq I$. Then, in both cases, $[P_{0},V]=\alpha\>V_{1}\otimes V_{2}$ for some $\alpha\in\mathbb{R}^{*}$. ∎
In the second Lemma we show that the commutators of commutators can be written as nested commutators by a simple re-ordering:
Lemma 2.
Let $A,B,C,D\in i\mathcal{P}_{N}$ and $[[A,B],[C,D]]\neq 0$. Then,
$[[A,B],[C,D]]\propto\begin{cases}ad_{D}ad_{C}ad_{A}(B)\>\>\mathrm{if}\>\>[AB,D]=0\\ ad_{C}ad_{D}ad_{A}(B)\>\>\mathrm{if}\>\>[AB,C]=0\end{cases}$ (38)
Proof.
We consider the binary notation as introduce in Appendix B.
$ad_{ad_{\bm{a}}(\bm{b})}(ad_{c}(\bm{d}))=\left[[\bm{a}^{\top}\Lambda\bm{b})](\bm{a}+\bm{b})^{\top}\Lambda[\bm{c}^{\top}\Lambda\bm{d}](\bm{c}+\bm{d})\right]\left([\bm{a}^{\top}\Lambda\bm{b}]\bm{a}+\bm{b}+[\bm{c}^{\top}\Lambda\bm{d}]\bm{c}+\bm{d}\right)\neq 0$
From this follows,
$[\bm{a}^{\top}\Lambda\bm{b}]=[\bm{c}^{\top}\Lambda\bm{d}]$ $=1$
$(\bm{a}+\bm{b})^{\top}\Lambda(\bm{c}+\bm{d})$ $=1\Leftrightarrow(\bm{a}+\bm{b})^{\top}\Lambda\bm{c}=1-(\bm{a}+\bm{b})^{\top}\Lambda\bm{d}.$
That means, there is two possibilities:
If $(\bm{a}+\bm{b})^{\top}\Lambda\bm{c}=1$, then, $(\bm{a}+\bm{b})^{\top}\Lambda\bm{d}=0$. Since $[\bm{c}^{\top}\Lambda\bm{d}]=1$, $(\bm{a}+\bm{b}+\bm{c})^{\top}\Lambda\bm{d}=1$ and the first case follows.
If $(\bm{a}+\bm{b})^{\top}\Lambda\bm{d}=1$, then, $(\bm{a}+\bm{b})^{\top}\Lambda\bm{c}=0$. Since $[\bm{c}^{\top}\Lambda\bm{d}]=1$, $(\bm{a}+\bm{b}+\bm{d})^{\top}\Lambda\bm{c}=1$ and the second case follows. ∎
Summarizing up to this point, Lemma 1 shows Proposition 2 for $\mathcal{A}=i\mathcal{P}_{k}$ and Lemma 2 shows that any commutator of commutators can be written as a nested commutator. That is, we can consider a sequence $P_{1},...,P_{r}\in i\mathcal{P}_{k}$ such that $V\propto ad_{P_{1}}\cdots ad_{P_{r}}(V^{\prime})$ is nonzero. Since $\langle\mathcal{A}\rangle_{\textrm{Lie}}=\mathfrak{su}(2^{k})$, we have that $P_{1},...,P_{r}\in\langle\mathcal{A}\rangle_{[\cdot,\cdot]}$. That is, there exist $A^{(i)}_{1},...,A^{(i)}_{s_{i}}\in\mathcal{A}$ such that,
$P_{i}=\beta_{i}\>ad_{A^{(i)}_{1}}\cdots ad_{A^{(i)}_{s_{i}-1}}(A^{(i)}_{s_{i}})$ (39)
for all $i=1,...,s$ and some $\beta\in\mathbb{R}^{*}$. Lemma 2 then shows that one can reorder the resulting commutators of commutators to yield a nested commutator of the desired form.
Appendix I Proof of 2
Using the notation established in Appendix B, let use define, for any $\mathcal{A}\subset i\mathcal{P}_{N}^{*}$, the following analogue of Equation 10:
$\langle f(\mathcal{A})\rangle_{[\cdot,\cdot]}=f(\mathcal{A})\bigcup_{r=1}^{\infty}\{ad_{\bm{v}_{i_{1}}}\cdots ad_{\bm{v}_{i_{r}}}(\bm{v})|(\bm{v}_{i_{1}},...,\bm{v}_{i_{r}},\bm{v})\in f(\mathcal{A})^{r+1}\}.$ (40)
As discussed in the main text, to prove 2, it suffices to consider sets $\mathcal{A}$ consisting of $2N$ independent elements. In particular, this means that $f(\mathcal{A})$ is a basis for $\mathbb{F}_{2}^{2N}$. Accordingly, proving 2 amounts to proving the following equivalent theorem:
Theorem 1.
Let $\widehat{\mathcal{A}}$ be a basis of $\mathbb{F}_{2}^{2N}$ where $N\geq 3$. Then
$\langle\widehat{\mathcal{A}}\rangle_{[\cdot,\cdot]}\neq\mathbb{F}_{2}^{2N}.$ (41)
Proof.
The proof proceeds by demonstrating that for any choice of $\widehat{\mathcal{A}}$, there always exists an element of $\mathbb{F}_{2}^{2N}$ that is not contained in $\langle\widehat{\mathcal{A}}\rangle_{[\cdot,\cdot]}$. We split this into two cases: (i) where $\widehat{\mathcal{A}}$ every element $\bm{v}\in\widehat{\mathcal{A}}$ is such that $\bm{v}^{\top}\Lambda\bm{w}=1$ for all $\bm{w}\in\widehat{\mathcal{A}}$ such that $\bm{w}\neq\bm{v}$, and (ii) where there exist distinct elements $\bm{v}$ and $\bm{w}$ in $\widehat{\mathcal{A}}$ such that $\bm{v}^{\top}\Lambda\bm{w}=0$. Case (i) is proven as 1 below while case (ii) is proven in 2 and 3. ∎
Proposition 1.
Let $\widehat{\mathcal{A}}$ be a basis for $\mathbb{F}_{2}^{2N}$ with $N\geq 3$, such that for all distinct $\bm{v},\bm{w}\in\widehat{\mathcal{A}}$, $\bm{v}^{\top}\Lambda\bm{w}=1$. Then for any three distinct $\bm{v},\bm{w},\bm{x}\in\widehat{\mathcal{A}}$, $\bm{v}+\bm{w}+\bm{x}\notin\langle\widehat{\mathcal{A}}\rangle_{[\cdot,\cdot]}$.
Proof.
For any sequence $\bm{u}_{1},...,\bm{u}_{r}\in\widehat{\mathcal{A}}$
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})=\bm{0}$ (42)
if $r\geq 3$ since
$ad_{\bm{u}_{r-2}}ad_{\bm{u}_{r-1}}(\bm{u}_{r})$ $=\bm{u}_{r-2}^{\top}\Lambda(\bm{u}_{r-1}+\bm{u}_{r})\left[\bm{u}_{r-2}+\bm{u}_{r-1}+\bm{u}_{r}\right]$ (43)
$=\bm{u}_{r-2}^{\top}\Lambda\bm{u}_{r-1}\left[\bm{u}_{r-2}+\bm{u}_{r-1}+\bm{u}_{r}\right]$
$\quad\quad+\bm{u}_{r-2}^{\top}\Lambda\bm{u}_{r}\left[\bm{u}_{r-2}+\bm{u}_{r-1}+\bm{u}_{r}\right]$ (44)
$=\bm{0}$ (45)
and since $ad_{\bm{u}_{l}}(\bm{0})=\bm{0}$. The result follows since there is no sequence $\bm{u}_{1},\bm{u}_{2}$ such that $ad_{\bm{u}_{1}}(\bm{u}_{2})=\bm{v}+\bm{w}+\bm{x}$. ∎
Proposition 2.
Let $\widehat{\mathcal{A}}$ be a basis for $\mathbb{F}_{2}^{2N}$ such that there exists distinct elements $\bm{v},\bm{w}\in\widehat{\mathcal{A}}$ satisfying $\bm{v}^{\top}\Lambda\bm{w}=0$. There does not exist a sequence $\bm{u}_{1},...,\bm{u}_{r}\in\widehat{\mathcal{A}}$ for odd $r$ such that
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})=\bm{v}+\bm{w}.$ (46)
Proof.
Assume that
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})=\bm{v}+\bm{w}.$ (47)
In order to not violate linear independence, we must have that
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})\neq\bm{0}$ (48)
and that $\bm{v}$ and $\bm{w}$ each appear in the sequence an odd number of times. We also know that, if $ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})\neq\bm{0}$ then
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})=\bm{u}_{1}+...+\bm{u}_{r}.$ (49)
It follows that
$\bm{u}_{1}+...+\bm{u}_{r}=\bm{v}+\bm{w}+\bm{u}_{l_{1}}+...+\bm{u}_{l_{t}}.$ (50)
If $r$ is odd then so is $t$, meaning that either $\bm{u}_{l_{1}}+...+\bm{u}_{l_{t}}\neq\bm{0}$, contradicting the assumption that $ad_{\bm{u}_{1}}...\bm{u}_{r-1}(\bm{u}_{r})=\bm{v}+\bm{w}$, or $\bm{u}_{l_{1}}+...+\bm{u}_{l_{t}}=\bm{0}$, contradicting the assumption of linear independence. ∎
Proposition 3.
Let $\widehat{\mathcal{A}}$ be a basis for $\mathbb{F}_{2}^{2N}$ that contains two distinct elements $\bm{v},\bm{w}$ such that $\bm{v}^{\top}\Lambda\bm{w}=0$. Then there does not exist a sequence $\bm{u}_{1},...,\bm{u}_{r}\in\widehat{\mathcal{A}}$ with $r=2k$, $k\in\mathbb{N}$ such that
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})=\bm{v}+\bm{w}.$ (51)
Proof.
The case where $r=2$ is trivial by the assumption. Suppose such a sequence $\bm{u}_{1},...,\bm{u}_{r}$ exists for $r=2k>2$. Since $\bm{v}+\bm{w}\neq\bm{0}$ by linear independence, this means that
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})\neq\bm{0}$ (52)
which is equivalent to
$\prod_{i=1}^{r-1}\left[\bm{u}_{i}^{\top}\Lambda(\bm{u}_{i+1}+...+\bm{u}_{r})\right].$ (53)
In particular, this implies that
$\sum_{i=1}^{r-1}\left[\bm{u}_{i}^{\top}\Lambda(\bm{u}_{i+1}+...+\bm{u}_{r})\right]=r-1\bmod 2=1.$ (54)
It is possible to rewrite the left-hand side of the above equation as
$\sum_{i=1}^{r-1}\sum_{j>i}\bm{u}_{i}^{\top}\Lambda\bm{u}_{j}.$ (55)
Since $ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})\neq\bm{0}$, we also have that
$ad_{\bm{u}_{1}}...ad_{\bm{u}_{r-1}}(\bm{u}_{r})=\bm{u}_{1}+...+\bm{u}_{r}.$ (56)
For $\bm{u}_{1}+...+\bm{u}_{r}$ to equal $\bm{v}+\bm{w}$, it must be the case that $\bm{v}$ and $\bm{w}$ appear in the sequence $\bm{u}_{1},...,\bm{u}_{r}$ an odd number of times and any other distinct element appears in the sequence an even number of times.
Let us consider Equation 55 for a sequence of the above type. Suppose without loss of generality that the first $\bm{v}$ appears before the first $\bm{w}$ in the sequence and let us write $(l_{(0,1)},l_{(0,2)})$ for the pair of corresponding indices, i.e. $\bm{u}_{l_{(0,1)}}=\bm{v}$ and $\bm{u}_{l_{(0,2)}}=\bm{w}$. The remaining $r-2$ elements of the sequence can be paired as follows: let $(l_{(i,1)},l_{(i,2)})$ for $i=1,...,k-1$ be a pair of labels such that $l_{(i,1)}$ is the label of the $i$th unpaired element of the sequence, $l_{(i,2)}>l_{(i,1)}$ has not yet been assigned to another pair, and $\bm{u}_{l_{(i,1)}}=\bm{u}_{l_{(i,2)}}$. Let us assume without loss of generality that $l_{(1,1)}<l_{(2,1)}<...<l_{(k-1,1)}$. By the assumptions on the sequence $\bm{u}_{1},...,\bm{u}_{r}$, it is possible to assign pairs in this way such that
$\{l_{(i,1)},l_{(i,2)}:i=0,...,k-1\}=\{1,...,r\}.$ (57)
Noting that it must be the case that the label corresponding to $r$ must be $l_{(i,2)}$ for some $i$, let us then define
$L_{1}$ :=$:=\{l_{(i,1)}:i=0,...,k-1\},$ (58)
$L_{2}$ :=$:=\{l_{(i,2)}:i=0,...,k-1\}\setminus\{r\}.$ (59)
We can then rewrite Equation 55 as
$\sum_{l_{(i,1)}\in L_{1}}\sum_{l_{(i,1)}<j}\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{u}_{j}+\sum_{l_{(i,2)}\in L_{2}}\sum_{l_{(i,2)}<j^{\prime}}\bm{u}_{l_{(i,2)}}^{\top}\Lambda\bm{u}_{j^{\prime}}.$ (60)
Using that $\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{u}_{l_{(i,2)}}=0$, either since $\bm{u}_{l_{(i,1)}}=\bm{u}_{l_{(i,2)}}$ for $i=1,...,k-1$ or since $\bm{v}^{\top}\Lambda\bm{w}=0$ by assumption, and using that addition is modulo $2$, the above expression can be written as:
$\sum_{\begin{subarray}{c}l_{(0,1)}<j,\\ j\neq l_{(0,2)}\end{subarray}}\bm{v}^{\top}\Lambda\bm{u}_{j}$ $+(1-\delta_{l_{(0,2),r}})\sum_{l_{(0,2)}<j^{\prime}}\bm{w}^{\top}\Lambda\bm{u}_{j^{\prime}}$
$+\sum_{\begin{subarray}{c}l_{(i,1)}\in L_{1},\\ i\neq 0\end{subarray}}\sum_{l_{(i,0)}<j^{\prime\prime}<l_{(i,2)}}\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{u}_{j^{\prime\prime}}.$ (61)
For every $j>l_{(0,1)}$ there are two possible cases: (i) $j=l_{(i,0)}$ for some $i$ meaning that $l_{(0,1)}<l_{(i,1)}<l_{(i,2)}$ or (ii) $j=l_{(i,2)}$ for some $i$ with $l_{(i,1)}<l_{(0,1)}$. In the first case both $\bm{v}^{\top}\Lambda\bm{u}_{l_{(i,1)}}$ and $\bm{v}^{\top}\Lambda\bm{u}_{l_{(i,2)}}$ appear in the first sum in the above expression, and since $\bm{u}_{l_{(i,1)}}=\bm{u}_{l_{(i,2)}}$, these terms cancel in the modulo two addition. For the second case, $\bm{v}^{\top}\Lambda\bm{u}_{l_{(i,2)}}$ appears in the first sum while $\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{v}$ appears in the third term in the above expression, so again there is cancellation (recall that $\bm{x}^{\top}\Lambda\bm{y}=\bm{y}^{\top}\Lambda\bm{x}$). Accordingly, the first term above fully cancels out. The exact same reasoning can be used to cancel the remaining terms, as follows. For the second term, either $l_{(0,2)}=r$ meaning that this term is zero anyway, otherwise any $j^{\prime}>l_{(0,2)}$ also satisfies the same two cases above, and the corresponding terms cancel out. For the remaining term, let us denote by $\widetilde{L}_{1}$ the set of labels $l_{(i,1)}$ that haven’t been cancelled out via the above method. For any $i,i^{\prime}$ such that $i<i^{\prime}$ and where $l_{(i,1)},l_{(i^{\prime},1)}\in\widetilde{L}_{1}$, we have the same two cases: either (i) $l_{(i,1)}<l_{(i^{\prime},1)}<l_{(i^{\prime},2)}<l_{(i,2)}$ in which case the terms $\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{u}_{l_{(i^{\prime},1)}}$ and $\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{u}_{l_{(i^{\prime},2)}}$ cancel, or (ii) $l_{(i,1)}<l_{(i^{\prime},1)}<l_{(i,2)}<l_{(i^{\prime},2)}$, in which case $\bm{u}_{l_{(i,1)}}^{\top}\Lambda\bm{u}_{l_{(i^{\prime},1)}}$ and $\bm{u}_{l_{(i^{\prime},2)}}^{\top}\Lambda\bm{u}_{l_{(i,2)}}$ cancel. (Note: there is a third case where $l_{(i,1)}<l_{(i,2)}<l_{(i^{\prime},1)}<l_{(i^{\prime},2)}$ but this case trivially contains no elements of the form $\bm{u}_{l_{(i,s)}}^{\top}\Lambda\bm{u}_{l_{(i^{\prime},s^{\prime})}}$, $s,s^{\prime}\in\{0,1\}$.)
We have thus shown that the left-hand side of Equation 54 must in fact be 00, forming the desired contradiction. ∎
Appendix J Proof of Theorem 3
For convenience, we restate the Theorem 3 here:
Theorem 1.
Consider a set of operators $\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ as in Theorem 1 such that $|\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}|=2N+1$. Let $V\otimes W\in i\mathcal{P}_{N}$. Then, PauliCompiler($V\otimes W$) of Algorithm 1 returns a sequence of operators $G_{1},...,G_{L}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ such that $V\otimes W\propto ad_{G_{1}}\cdots ad_{G_{L-1}}(G_{L})$ and $L=\mathcal{O}(N)$.
Proof.
Algorithms 1 and 2 have been reprinted in Figs. 1 and 2, respectively. We refer to the line numbering there in the following.
We have proven through Lemmas 1 and 1, that PauliCompiler deterministically returns such a sequence $G_{1},...,G_{L}\in\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}$ in finite time.
Now, let us prove $L=\mathcal{O}(N)$. With the exception of the case $W=I$ in line 2 of Fig. 1, all returned sequence rely on the output of SubsystemCompiler. The case $W=I$ returns a sequence that is constant in length since $s$ is constant. Therefore, let us consider the length of a sequence $\mathcal{G}$ returned by SubsystemCompiler in Fig. 2. It follows immediately from the proof of Lemma 1 in Appendix G that the worst case length of the returned sequence $\mathcal{G}$ is $4(N-k)$. To see this consider $|\mathcal{B}^{\prime}|=2(N-k)$ and $|\mathcal{A}^{\prime}|=2k+1$. Then, the worst case length for an optimal choice of a product $\prod_{i=1,...,r}B_{i}$ to realize another Pauli operator proportional to $W\in i\mathcal{P}_{N-k}$ is $r=N-k$. This is the case for $W\propto\prod_{B\in\mathcal{B}}B$. Since the set $\mathcal{B}$ only contains independent elements, any other operator in $i\mathcal{P}_{N-k}$ can be written as a product with fewer factors in $\mathcal{B}$. This is obvious when we consider that $\mathcal{B}$ forms a basis of $\mathbb{F}_{2}^{2N}$ as discussed in Appendix B. In the worst case, Algorithm 2 then adds, for each factor, three additional elements (in lines 9 and 13, respectively). Hence, $|\mathcal{G}|=4(N-k)$ in the worst case.
The longest worst case sequence of PauliCompiler in Fig. 1 is returned in line 11 if both sequences $\mathcal{G}^{\prime}$ in line 7 and $\mathcal{G}^{\prime\prime}$ in line 8 have length $4(N-k)$, respectively. The worst case length of the returned sequence is therefore, $L=8(N-k)+s$. Since $s$ is chosen to be a constant in line 9, $L=\mathcal{O}(N)$ in the worst case. ∎
Note that we consider the case $|\mathcal{A}^{\prime}\cup\mathcal{B}^{\prime}|=2N+1$ because otherwise, we could choose, for example, $\mathcal{B}=i\mathcal{P}_{N-k}$ such that any sequence returned by SubsystemCompiler can be chosen to be of size 1.